### v1

# FedALT: Federated Fine-Tuning through Adaptive Local Training with Rest-of-the-World LoRA

Abstract
Fine-tuning large language models (LLMs) in federated settings enables privacy-preserving adaptation but suffers from cross-client interference due to model aggregation. Existing federated LoRA fine-tuning methods, primarily based on FedAvg, struggle with data heterogeneity, leading to harmful cross-client interference and suboptimal personalization. In this work, we propose FedALT, a novel personalized federated LoRA fine-tuning algorithm that fundamentally departs from FedAvg. Instead of using an aggregated model to initialize local training, each client continues training its individual LoRA while incorporating shared knowledge through a separate Rest-of-the-World (RoTW) LoRA component. To effectively balance local adaptation and global information, FedALT introduces an adaptive mixer that dynamically learns input-specific weightings between the individual and RoTW LoRA components using the Mixture-of-Experts (MoE) principle. Through extensive experiments on NLP benchmarks, we demonstrate that FedALT significantly outperforms state-of-the-art personalized federated LoRA fine-tuning methods, achieving superior local adaptation without sacrificing computational efficiency.
Machine Learning, ICML
1 Introduction
Figure 1: Illustration of FedALT. Instead of directly aggregating local LoRA modules from each client using FedAvg, FedALT introduces a frozen Rest-of-World LoRA component to transmit shared global knowledge while preserving client-specific adaptations through Individual LoRA modules. The adaptive mixer dynamically combines the RoTW LoRA and Individual LoRA. See details in Section 3.
Large language models (LLMs) (Kenton & Toutanova, 2019; Brown et al., 2020; Raffel et al., 2020; Touvron et al., 2023a; Zhou et al., 2024; Zeng et al., 2022; Touvron et al., 2023b) have demonstrated exceptional capabilities in language understanding and generation, enabling a wide range of applications. Adapting pretrained LLMs to specialized domains or further enhancing their performance requires fine-tuning with task-specific data (Minaee et al., 2024). However, full fine-tuning, which involves updating all model parameters, is computationally expensive and often impractical for real-world deployment. To address this limitation, several parameter-efficient fine-tuning (PEFT) (Ding et al., 2023; Fu et al., 2023; Han et al., 2024; Liu et al., 2022; Li & Liang, 2021; Lester et al., 2021) methods have been proposed, with low-rank adaptation (LoRA) (Hu et al., 2021) being one of the most widely used. LoRA reduces the number of trainable parameters by integrating low-rank matrices into the model, significantly reducing computational costs.
Fine-tuning LLMs requires substantial data to adapt to downstream tasks, yet this data is often distributed across multiple institutions where privacy regulations and security concerns prohibit direct sharing. Federated learning (FL) (McMahan et al., 2017) enables collaborative fine-tuning without exposing raw data, making it an essential framework for privacy-preserving adaptation. However, most existing federated LoRA fine-tuning approaches (Zhang et al., 2024; Sun et al., 2024; Wang et al., 2024d; Guo et al., 2024; Bian et al., 2024b) assume a single global model, overlooking the highly heterogeneous data distributions among clients. A one-size-fits-all model often results in suboptimal performance and fairness issues, especially in real-world scenarios where client data varies significantly in size, domain, and distribution. For example, in healthcare (Kim et al., 2024), hospitals may require fine-tuned models tailored to specific patient demographics or disease profiles. Without personalization, federated fine-tuning fails to capture such variations, leading to poor local adaptation and inefficient resource utilization. Despite the critical need for personalized solutions, personalized federated LoRA fine-tuning (Yang et al., 2024a; Anonymous, 2025; Qi et al., 2024) remains largely underexplored.
Federated Averaging (FedAvg) (McMahan et al., 2017) is perhaps the most widely used FL framework, serving as the foundation for numerous follow-up works that introduce various optimizations. Its core principle is to aggregate locally trained client models into a global model, which then initializes the next round of local training. Most personalized FL algorithms (Yang et al., 2024a; Qi et al., 2024; Tan et al., 2022; Deng et al., 2020; Collins et al., 2021; Fallah et al., 2020), despite their implementation differences, build upon the FedAvg framework by incorporating personalization mechanisms—and personalized federated LoRA fine-tuning follows the same paradigm. For example, the recently proposed FedDPA (Yang et al., 2024a) algorithm personalizes LoRA by introducing both global and local LoRA components. The global LoRA component is trained using FedAvg, while the local LoRA component is trained independently on each client’s private data. This local adaptation can occur either after the global LoRA stabilizes or alternatedly with global LoRA training, in which case the final output is computed as a weighted combination of both components using a pre-defined weighting factor. However, while personalization improves the average performance across clients, some clients may experience degraded performance, even falling behind simple local fine-tuning—that is, not participating in FL at all. This issue, also observed in our motivational study in Section 2.3, raises a critical question: Is FedAvg truly the most effective framework for developing personalized federated LoRA fine-tuning methods in heterogeneous settings?
In this paper, we propose a novel personalized federated LoRA fine-tuning algorithm that departs fundamentally from the FedAvg paradigm. Our method is motivated by identifying two key limitations of FedAvg-based approaches, such as FedDPA, in handling FL challenges under heterogeneous client settings. First, FedAvg-based approaches suffer from harmful cross-client interference due to their reliance on a global model aggregated from client models. This aggregation step, intended to capture shared knowledge, often disrupts individual client adaptations when data distributions are substantially different. As a result, any improvements achieved during local fine-tuning can be negated by aggregation, making it ineffective for clients whose distributions deviate significantly from the global model. For such clients, initializing the next training round with the aggregated model may offer little to no benefit—or even degrade performance. Second, FedAvg-based approaches lack an effective mechanism to balance local and global information in personalization. For example, FedDPA combines global and local LoRA components using a pre-defined weighting factor during training. However, improper weighting can significantly harm personalized fine-tuning, particularly when the global model is assigned excessive weight. This issue is especially problematic for clients where the global model is poorly suited, leading to suboptimal performance despite the presence of a local adaptation mechanism.
Overall, the FedAvg paradigm structures the personalization mechanisms around the aggregated model, treating local training as a supplement to the global model. In contrast, our approach emphasizes the role of local training in heterogeneous settings, focusing on personalization by incorporating useful global information directly into local training. Specifically, we eliminate the key FedAvg step of using the aggregated model to initialize each local training round. Instead, clients continue training on their own local models from the previous round. Information from other clients’ models is aggregated into a “rest-of-the-world” (RoTW) LoRA component, which supports local training but remains frozen during each round. Furthermore, to effectively leverage the RoTW component, we introduce an “adaptive mixer” that dynamically learns the optimal weighting between the individual LoRA and RoTW LoRA components using the Mixture-of-Experts (MoE) principle (Jordan & Jacobs, 1994). This adaptive weighting enables the model to selectively integrate relevant global knowledge while maintaining local model adaptation. We name our approach FedALT, short for Federated Fine-tuning through Adaptive Local Training with Rest-of-the-World LoRA.
We highlight that FedALT is particularly well-suited for federated LoRA fine-tuning due to its low computational complexity and minimal memory requirements. Although a separate RoTW LoRA is maintained alongside each client’s individual LoRA component, the RoTW LoRA is not trained during each round. Furthermore, LoRA modules are significantly smaller than the full model, meaning the added memory overhead is negligible. Through comprehensive experiments on heterogeneous FL benchmarks across various NLP tasks, we demonstrate that FedALT outperforms state-of-the-art approaches by a significant margin.
2 Preliminaries and Motivations
2.1 LoRA
Consider a pre-trained model with parameters $\mathbf{W}_{0}\in\mathbb{R}^{l\times d}$, where $\mathbf{W}_{0}$ represents the fixed parameters of the model, and $\mathbf{\Delta W}\in\mathbb{R}^{l\times d}$ denotes the trainable update matrix applied during fine-tuning. Here, $d$ is the input dimension and $l$ is the output dimension. Instead of updating all elements in $\mathbf{\Delta W}$, LoRA decomposes $\mathbf{\Delta W}$ into two low-rank matrices $\mathbf{A}\in\mathbb{R}^{r\times d}$ and $\mathbf{B}\in\mathbb{R}^{l\times r}$, where $r\ll\min(d,l)$. This decomposition allows the fine-tuning process to focus on the significantly smaller low-rank matrices $\mathbf{A}$ and $\mathbf{B}$ instead of the full matrix $\mathbf{\Delta W}$. Consequently, the total number of trainable parameters is reduced from $d\times l$ to $r\times(d+l)$. The model parameters after fine-tuning are given by:
$\mathbf{W}=\mathbf{W}_{0}+\mathbf{\Delta W}=\mathbf{W}_{0}+\mathbf{B}\mathbf{A},$ (1)
where $\mathbf{A}$ is typically initialized with random Gaussian values, while $\mathbf{B}$ is initialized to zero, ensuring a stable start to the fine-tuning process.
2.2 Heterogeneous Federated Fine-Tuning
Assume there are $K$ institutions (clients), each possessing a distinct local training dataset $\mathcal{D}_{k}=\{X_{k},Y_{k}\}$, where $k$ indexes a client. In the context of LLM fine-tuning, it is common for clients to have datasets corresponding to entirely different tasks (e.g., one client may have data for text summarization, while another client has data for sentiment analysis), rather than simply exhibiting statistical differences as in conventional FL.
In this scenario, each client $k$ aims to leverage FL to extract useful information while fine-tuning a model tailored to its specific task. The goal is for each client to learn an individualized fine-tuned model $\textbf{W}_{k}$ that best fits its local dataset. The objective of the training process can be formulated as:
$\min_{\mathcal{W}}\frac{1}{K}\sum_{k=1}^{K}L(X_{k},Y_{k},\mathbf{W}_{k}),$ (2)
where $L(\cdot)$ represents the loss function, and $\mathcal{W}$ denotes the set of personalized fine-tuned models $\{\mathbf{W}_{k}\}_{k=1}^{K}$.
2.3 Motivational Study
Federated fine-tuning operates in highly heterogeneous environments where client datasets can differ significantly. To understand the effects of such heterogeneity, we conducted a motivational study that highlights both the benefits and challenges of applying FL in these settings.
Figure 2: Motivational study results shows that while FedIT underperforms Local Only on some clients due to harmful cross-client interference. FedIT-split, using multiple smaller LoRAs, fails to mitigate this interference, highlighting the limitations of the FedAvg paradigm. See details in Section 2.3.
Fact 1: FL Can Transfer Useful Knowledge but Also Introduces Harmful Interference.
In FL settings, the data collected by each client can be highly diverse (Huang et al., 2022; Mendieta et al., 2022; Wang et al., 2024b). We conducted a motivational experiment using the Flan dataset (Chung et al., 2024), selecting 8 tasks and creating 8 clients, each specializing in a single task. To demonstrate potential interference in federated fine-tuning, we compared two baseline methods. The first method, FedIT (Zhang et al., 2024), combines FedAvg with LoRA fine-tuning. The second method is Local Only LoRA fine-tuning, where each client fine-tunes its LoRA locally without any communication or information sharing between clients. The results, shown in Figure 2, support our first observation. On one hand, federated fine-tuning using FedIT achieves better performance than local fine-tuning on tasks Coreference Resolution and Paraphrase Detection, confirming that FL can enhance learning by enabling clients to benefit from shared knowledge. On the other hand, for tasks Commonsense Reasoning and Text Classification, FedIT underperforms compared to individual local fine-tuning. This dual outcome underscores both the promise and the pitfalls of federated fine-tuning in heterogeneous settings.
Fact 2: FedAvg Paradigm Struggles to Mitigate Interference
Having established the existence of harmful interference, we next explored potential mitigation strategies. In centralized LoRA fine-tuning, HydraLoRA (Tian et al., 2024) found that splitting a single large LoRA into multiple smaller LoRAs reduces cross-domain interference. However, FL introduces additional challenges, as data domains are distributed across clients. To test whether multiple smaller LoRAs could mitigate interference in FL, we implemented FedAvg with multiple parallel LoRAs (i.e. FedIT-split). However, as shown in Figure 2, its performance remained comparable to FedIT (FedAvg with a single large LoRA), indicating that interference persisted. This result is expected: in centralized settings, interference occurs within a shared dataset and can be mitigated by partitioning LoRAs. In FL, interference arises primarily during server aggregation, and simply introducing multiple LoRAs does not prevent the aggregated model from disrupting local adaptations. Consequently, direct adaptations of centralized interference-mitigation strategies fail to address the core issue in FL.
3 Proposed Method: FedALT
Our motivational studies reveal that harmful interference in federated learning arises from server aggregation and the use of the aggregated model to initialize subsequent training rounds. In the FedAvg paradigm, a global LoRA serves as a shared initialization point, allowing clients to fine-tune locally on their respective datasets. After fine-tuning, the updated LoRAs are averaged on the server and redistributed to clients as the new initialization. However, when client data distributions differ significantly, aggregation disrupts individual adaptations, often negating improvements from local training and hindering personalization.
Can cross-client interference be mitigated? The simplest and most direct solution is local fine-tuning using only a client’s own dataset. This approach eliminates interference entirely but fails to leverage shared knowledge from other clients. To address this trade-off, we propose a novel personalized federated LoRA fine-tuning algorithm that fundamentally diverges from the FedAvg paradigm, prioritizing local training while selectively integrating useful global information.
Our method introduces two key innovations: 1. Each client maintains two LoRA modules: an Individual LoRA, which captures locally learned information and is updated exclusively during the client’s fine-tuning, and a Rest-of-the-World (RoTW) LoRA, which aggregates information from all other clients. Crucially, the RoTW LoRA remains fixed during local fine-tuning, ensuring that aggregation does not interfere with a client’s adaptation. 2. An adaptive mixer dynamically learns the optimal input-specific weighting between the Individual LoRA and RoTW LoRA components, leveraging the Mixture-of-Experts (MoE) principle to balance personalization and global knowledge integration.
In the following of this section, we detail the design and functionality of these components.
3.1 Individual and Rest-of-the-World (RoTW) LoRA
Each client $k$ maintains two types of LoRAs: an Individual LoRA, denoted by $\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$, and a Rest-of-the-World (RoTW) LoRA, denoted by $\mathbf{A}^{R}_{k}/\mathbf{B}^{R}_{k}$. The Individual LoRA captures information obtained locally by the client and is isolated from the information of other clients. In contrast, the RoTW LoRA represents the average of the Individual LoRAs from all other clients and remains frozen during the client’s local fine-tuning. Specifically, at the end of each round, client $k$ updates its Individual LoRA $\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$ through local training. The RoTW LoRA for client $k$ is then computed as:
$\mathbf{A}^{R}_{k}=\frac{1}{K-1}\sum_{m\neq k}\mathbf{A}^{L}_{m},\quad\mathbf{B}^{R}_{k}=\frac{1}{K-1}\sum_{m\neq k}\mathbf{B}^{L}_{m},$ (3)
After clients upload their local LoRAs to the server, the server computes the RoTW LoRA for each client individually and distributes it back to them. Alternatively, the server can calculate the global average of all clients’ individual LoRAs and distribute this average back to the clients. Each client then determines its RoTW LoRA by subtracting its individual LoRA from the global average LoRA.
Now, consider the scenario where a weight $\frac{1}{K}$ is assigned to the local LoRA and $1-\frac{1}{K}$ to the RoTW LoRA when combining their outputs. In this case, the combined LoRA is mathematically equivalent to the averaged LoRA. This equivalence arises naturally from the definitions of the RoTW and averaged LoRAs. If these two LoRAs are trained simultaneously on local datasets, the approach essentially reduces to the standard FedAvg (or FedIT) algorithm, as all clients would end up with the same combined LoRA.
However, it is crucial to emphasize that the RoTW LoRA is not further updated on clients using their local datasets. Instead, it is only updated through averaging at the end of each learning round. Training the RoTW LoRA on local datasets is unnecessary and potentially counterproductive, as subsequent averaging could cancel out any improvements. Furthermore, skipping local training for the RoTW LoRA reduces the computational workload on clients by half.
3.2 Dynamic Mixture-of-Experts Mechanism
A natural question arises: why not simply add the RoTW LoRA directly to the pre-trained model, i.e., $\mathbf{W}_{0}\leftarrow\mathbf{W}_{0}+\mathbf{B}^{R}_{k}\mathbf{A}^{R}_{k}$, since the RoTW LoRA remains fixed during local training? This approach would allow each client to maintain only a single Individual LoRA, saving memory space. However, there are two key reasons why this is not desirable:
  1. 1.
    Potential Contamination of the Pre-trained Model: If the RoTW LoRA underperforms, adding it directly to the pre-trained model risks contaminating it, making further corrections difficult.
  2. 2.
    Loss of Flexibility: Directly incorporating the RoTW LoRA enforces a fixed weight, preventing dynamic adjustment between the Individual LoRA and shared information for different inputs. This lack of flexibility is problematic because different inputs may benefit variably from the local model versus the globally averaged model. A “one-size-fits-all” approach often fails to perform optimally across diverse inputs.
To address these challenges, we propose a dynamic weighting mechanism for combining the Individual LoRA and the RoTW LoRA, enabling input-specific flexibility. Specifically, we introduce a mixer to dynamically adjust the contributions of the two LoRAs for each input. During the forward pass of each local epoch, the model’s output is computed as:
$y=\mathbf{W}_{0}x+\alpha_{k}(x)\mathbf{B}^{L}_{k}\mathbf{A}^{L}_{k}x+\big{(}1-\alpha_{k}(x)\big{)}\mathbf{B}^{R}_{k}\mathbf{A}^{R}_{k}x,$
where $\alpha_{k}(x)$ determines for client $k$ the contribution weight of the Individual LoRA, and $1-\alpha_{k}(x)$ determines the contribution of the RoTW LoRA.
Given the complexity of modern LLMs, where LoRA modules are inserted at multiple layers, learning $\alpha(x)$ is non-trivial. Inspired by (Jordan & Jacobs, 1994), we adopt a Mixture-of-Experts (MoE) approach to compute $\alpha(x)$. Specifically, we introduce a mixer $\mathbf{G}_{k}\in\mathbb{R}^{2\times d}$ which is a dense layer with trainable weights (a transformation matrix), followed by a softmax function, to learn the weighting:
$\alpha(x),\,1-\alpha(x)=\text{softmax}(\mathbf{G}_{k}x).$ (4)
Importantly, the MoE mixers are personalized for each client, tailored to their specific domain tasks, and are not averaged across clients. This personalization ensures that the weight adjustments reflect each client’s unique data distribution and learning objectives.
3.3 Algorithm Workflow
By combining the two LoRA modules (Individual LoRA and RoTW LoRA) with dynamic weights, our proposed method achieves a balance between leveraging useful information from other clients and isolating harmful interference. To better illustrate the FedALT training process, we provide an overview of the server and client operations during fine-tuning at round $t$.
Server Side. After local fine-tuning in round $t$, the server receives each client $k$’s Individual LoRA modules ($\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$). Upon receiving all clients’ Individual LoRA modules, the server updates and computes the RoTW LoRA modules for each client $k$ based on Equation 3. After calculating the RoTW LoRA modules for client $k$, the server broadcasts these updated RoTW LoRA modules to the respective client.
Client Side. Each client $k$ replaces its current RoTW LoRA modules with the new RoTW LoRA modules ($\mathbf{A}^{R}_{k}/\mathbf{B}^{R}_{k}$) received from the server. It then begins local fine-tuning for round $t+1$ using its local dataset. During the local fine-tuning process, the client updates the parameters in its Individual LoRA modules ($\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$) and the mixer $\mathbf{G}_{k}$, while the RoTW LoRA modules ($\mathbf{A}^{R}_{k}/\mathbf{B}^{R}_{k}$) and the pre-trained model ($\mathbf{W}_{0}$) remain fixed. Once the local fine-tuning is complete, the updated Individual LoRA is uploaded to the server, while $\mathbf{G}_{k}$ stays local to the client.
3.4 Discussion: Why a Single RoTW LoRA?
FedALT prioritizes local training while selectively incorporating useful information from other clients via a dynamic weighting mechanism. Ideally, each client could tailor this incorporation by maintaining separate LoRA components for every other client. In this approach, client $k$ would receive the local LoRA components $\mathbf{A}^{L}_{j}/\mathbf{B}^{L}_{j}$ from all other clients $j\neq k$, freeze them during local training, and learn separate weights for each $j$:
$y=\mathbf{W}_{0}x+\alpha_{k,k}(x)\mathbf{B}^{L}_{k}\mathbf{A}^{L}_{k}x+\sum_{j\neq k}\alpha_{k,j}(x)\mathbf{B}^{L}_{j}\mathbf{A}^{L}_{j}x$
where $\sum_{j}\alpha_{k,j}(x)=1$. This would allow client $k$ to capture the distinct influence of each other client’s LoRA.
However, this approach presents significant drawbacks. First, each client would need to store and communicate a LoRA component for every other client, leading to memory and bandwidth overhead that scales linearly with the number of clients $K$, making it impractical for large-scale settings. Second, training the mixer becomes increasingly complex with more LoRA components, requiring additional parameters and longer convergence times. To address these challenges, we adopt a single RoTW LoRA, which aggregates knowledge from other clients into a single component. While this reduces computational and communication costs, it may limit personalization. Future work could explore clustering-based approaches to maintain a small number of RoTW LoRA components, capturing finer-grained client relationships while remaining scalable.
4 Experiments
In this section, we evaluate and compare the performance of our proposed method, FedALT, with other baseline methods across two federated learning dataset settings. All experiments are conducted on four NVIDIA A100 GPUs. Detailed training implementations are provided in the Appendix B. The code and datasets used for the experiments will be made publicly available upon acceptance of the paper.
Datasets. We construct our training datasets from Flan (Chung et al., 2024), a comprehensive collection of over 60 NLP datasets covering more than 12 diverse NLP tasks. For our experiments, we consider two federated dataset settings. In the first setting, we simulate extreme client heterogeneity by assuming 8 clients, each having a completely different NLP task dataset. In the second setting, we assume 8 clients with datasets spanning a total of 6 NLP tasks. Unlike the first setting, some clients in the second setting share the same NLP task, even though their datasets come from different sources. Detailed descriptions of the datasets, including task distributions and client assignments, are provided in the Appendix A.
Pre-trained Model and Baselines. We utilize LLaMA2-7B (Touvron et al., 2023b) as the pre-trained model for all baseline and proposed fine-tuning methods. To evaluate the performance of our proposed method, we compare it with the state-of-the-art methods in federated fine-tuning with LoRA, including FedIT (Zhang et al., 2024), FFA-LoRA (Sun et al., 2024), FedSA (Guo et al., 2024). Additionally, we compare it with state-of-the-art methods focused on personalized federated fine-tuning with LoRA, such as FedDPA (Yang et al., 2024a), PF2LoRA (Anonymous, 2025), and FDLoRA (Qi et al., 2024). We also include the Local Only baseline, where clients fine-tune their models independently without federated collaboration. Detailed descriptions of these baseline methods are provided in the Appendix C.
Methods Dataset 1
Commonsense Coreference Natural Language Paraphrase Reading Sentiment Struct to Text Average
Reasoning Resolution Inference Detection Comprehension Analysis Text Classification
Local Only 73.83 74.62 58.73 71.00 55.82 46.56 55.14 67.18 62.81
FedIT 72.82 77.14 59.53 72.33 53.35 46.76 49.23 66.39 62.22
FFA-LoRA 66.96 68.48 50.41 69.79 49.69 44.44 43.29 61.15 55.54
FedSA 72.68 78.24 64.46 76.33 54.32 42.71 53.57 65.47 63.47
FedDPA 74.81 81.88 62.92 76.33 55.91 47.86 52.02 65.42 64.64
PF2LoRA 74.13 77.55 64.17 78.33 55.36 48.65 53.44 63.90 64.44
FDLoRA 76.29 75.60 66.12 75.67 57.39 49.85 52.85 67.59 65.17
FedALT 76.12 83.04 67.73 77.80 59.41 51.57 53.12 71.60 67.55
Table 1: Performance comparison with baseline methods in the first federated dataset setting. The average column represents the overall performance across these tasks. See details in Section 4.1.
Methods Dataset 2
Struct to Struct to Coreference Natural Language Natural Language Paraphrase Text Sentiment Average
Text 1 Text 2 Resolution Inference 1 Inference 2 Detection Classification Analysis
Local Only 52.98 57.35 74.08 66.10 69.02 78.67 64.97 70.16 66.67
FedIT 48.30 52.81 76.51 65.61 69.43 71.56 60.08 71.90 64.53
FFA-LoRA 50.64 51.70 68.09 60.99 64.59 70.00 61.47 69.68 62.15
FedSA 52.22 56.59 78.11 71.41 69.13 80.00 65.15 66.92 67.44
FedDPA 51.99 55.62 78.60 66.82 68.49 78.67 66.09 73.60 67.49
PF2LoRA 50.70 60.19 77.80 68.07 69.21 72.67 71.10 71.07 67.60
FDLoRA 52.71 58.37 74.74 69.62 66.27 71.00 65.99 73.80 66.56
FedALT 51.73 60.72 81.91 74.51 71.12 80.33 70.28 73.22 70.48
Table 2: Performance comparison with baseline methods in the second federated dataset setting. The average column represents the overall performance across these tasks. See details in Section 4.1.
4.1 Experiments Results
Overall Performance Comparisons. We compare the performance of the proposed method, FedALT, with other baseline methods based on each client’s performance on its specific test dataset, aligned with its training dataset. As shown in Table 1, which reports the results for the first federated setting, our proposed method achieves the best overall performance compared to all baselines, with the average score exceeding that of the best-performing baseline, FDLoRA, by 2.13. FedALT consistently delivers superior performance across most tasks, outperforming the baselines by at least 3.27 on the Natural Language Inference task and 3.50 on the Reading Comprehension task. We find that on the Struct-to-Text, the Local Only method achieves the highest performance. This is likely due to the unique nature of this task, making it difficult to benefit from shared global knowledge. Furthermore, we observe that FedIT underperforms compared to Local Only on four client tasks, indicating the presence of harmful interference during aggregation. Personalized federated LoRA fine-tuning methods (i.e., FedDPA, PF2LoRA, FDLoRA) also fall short compared to FedALT, likely because they fail to effectively isolate harmful interference during the server aggregation step. Despite their designs for local/global LoRA interaction, these methods update LoRA modules using aggregated global information, which may carry conflicting information from other clients.
To further validate the effectiveness of FedALT, Table 2 presents the results for the second federated setting, involving a different mix of NLP tasks, clients, and task types. The results align with those in the first setting, with FedALT outperforming all baseline methods. Notably, FedALT achieves an average score of 70.48, surpassing the baseline method by at least 2.88. This consistent performance across diverse settings highlights the robustness and adaptability of FedALT, demonstrating its ability to handle heterogeneous data distributions effectively.
Additionally, in Table 3, we report the number of trainable parameters and the parameters required for inference for both the baseline methods and our proposed method. Due to space limitations, we present FedIT as an example, as the other federated learning-based methods have identical trainable and inference parameters. All methods exhibit a similar number of trainable parameters, except for FFA-LoRA, which has the fewest trainable parameters since it only updates the LoRA $\mathbf{B}$ module. However, despite its reduced parameter count, FFA-LoRA is significantly outperformed by our proposed method in overall performance. Regarding the parameters required for inference, our method involves an additional computation for the RoTW LoRA. However, this parameter overhead is negligible compared to the size of the large pre-trained LLM.
Methods Trainable Parameters Additional Inference Workload
FedIT 0.0622% 0.0622%
FFA-LoRA 0.0311% 0.0311%
FedDPA 0.0622% 0.1245%
FedALT 0.0661% 0.1283%
Table 3: Comparison of trainable and inference parameters. Note that, except for FFA-LoRA and FedDPA, all other baselines share the same parameter count as FedIT. Our method employs a comparable number of LoRA parameters to other baselines, which is negligible compared to the pre-trained LLM. See details in Section 4.1.
Impact of Decoupling LoRA Training. In our proposed method, only the Individual LoRA modules are updated locally during client fine-tuning, while the RoTW LoRA remains frozen and is updated solely through server-side aggregation. This design aims to mitigate harmful cross-client interference by preventing local updates to the shared global component. To validate the effectiveness of this decoupled training approach, we compare FedALT against two alternative methods. The first comparison method, referred to as RoTW-Update (rank = 4), fine-tunes both the Individual LoRA and RoTW LoRA locally, with each LoRA having a rank of 4 to maintain an equivalent total number of trainable parameters. The second comparison method, RoTW-Update (rank = 8), also fine-tunes both LoRA modules locally but assigns a rank of 8 to each, resulting in a higher total parameter count. The results, shown in Table 4, reveal that while the second method outperforms the first due to its larger parameter count, both alternative methods underperform compared to our proposed approach. Notably, even with more trainable parameters, the second method fails to match the performance of FedALT.
Methods Same Trainable Param. Same Inference Param. Avg. Performance
RoTW-Update (rank = 4) $\checkmark$ $\times$ 63.59
RoTW-Update (rank = 8) $\times$ $\checkmark$ 65.75
FedALT 67.55
Table 4: Impact of decoupling LoRA training. Updating the RoTW LoRA component locally leads to harmful cross-client interference. See details in Section 4.1.
Figure 3: Impact of Dynamic Weighting via MOE. The proposed dynamic personalized mixer mechanism surpasses both fixed-weight approaches and shared averaged mixers across clients. See Section 4.1 for details.
Impact of Dynamic Weighting via MOE. To enable input-specific flexibility, our proposed method adopts a MOE mechanism to dynamically adjust the contributions of the Individual LoRA and RoTW LoRA. We conduct experiments to evaluate the efficiency of this dynamic weighting by comparing it to two fixed weighting strategies. In the first fixed strategy, we set the contribution weight to a fixed value of $\alpha=0.5$, meaning that the Individual LoRA and RoTW LoRA contribute equally. In the second fixed strategy, we set the contribution weight to $\alpha=1/K$, where $K$ is the number of clients. This configuration effectively reduces the method to using the averaged LoRA, relying solely on each client’s last-round LoRA modules for the next round’s initialization. The results in Figure 3 validate that our proposed dynamic weighting mechanism outperforms both fixed strategies, demonstrating its ability to enhance performance. Additionally, our dynamic weighting is implemented through the personal $\mathbf{G}_{k}$ mixer, which is maintained locally by each client and is not aggregated during the training process. To further investigate this design choice, we conducted experiments where $\mathbf{G}_{k}$ is averaged across clients. The results confirm that retaining a personalized $\mathbf{G}_{k}$ for each client yields optimal performance, emphasizing the importance of input-specific and client-specific adjustments.
5 Related Work
5.1 Parameter-Efficient Fine-Tuning
As the size of LLMs continues to grow, full parameter fine-tuning has become increasingly computationally and storage-intensive (Hadi et al., 2023). To address this issue, Parameter-Efficient Fine-Tuning (PEFT) methods (Ding et al., 2023; Fu et al., 2023; Han et al., 2024; Liu et al., 2022; Li & Liang, 2021; Lester et al., 2021) have been developed to significantly reduce the number of trainable parameters. PEFT techniques introduce a limited set of additional trainable parameters to enhance model performance while keeping the majority of pre-trained parameters frozen. Among all existing PEFT methods, the most prominent approach is LoRA (Hu et al., 2021), which employs low-rank matrices to approximate the pre-trained weight matrix, updating only the low-rank components.
LoRA has gained significant attention recently, becoming a standard method for adapting LLMs like Llama under limited computational resources (Lermen et al., 2023). Several works have been proposed to improve LoRA (Tian et al., 2024; Lermen et al., 2023; Wang et al., 2023; Sheng et al., 2023; Liu et al., 2023). For instance, a recent development, Hydra-LoRA (Tian et al., 2024), employs an asymmetric structure with a shared matrix for all samples and distinct matrices for each intrinsic component, thereby enhancing domain adaptation. While these advancements primarily focus on centralized learning settings, our work extends the application of PEFT methods to federated LLM fine-tuning, addressing the unique challenges posed by distributed and heterogeneous client data.
5.2 Federated Learning
The foundational work in federated learning, FedAvg (McMahan et al., 2017), addresses privacy and communication efficiency by aggregating local model parameters to train a shared global model. Since then, numerous studies (Wang et al., 2024a, c; Yang et al., 2024b; Bian et al., 2024a) have focused on tackling various challenges within FL settings. One of the key challenges is data heterogeneity, which makes it difficult to train a single shared global model that performs well across all clients. To address this, Personalized Federated Learning(PFL) (Tan et al., 2022; Deng et al., 2020; Collins et al., 2021; Fallah et al., 2020) has emerged as an approach to adapt the global model to the specific needs of each client in the federation. Most PFL research has focused on addressing statistical differences in data distributions (e.g., label distribution skew) and has been applied primarily to smaller, standard models. In contrast, our work addresses the unique challenges posed by heterogeneous clients with diverse NLP tasks by leveraging LLM fine-tuning within the FL framework.
5.3 Federated Fine-Tuning
Several studies (Bai et al., 2024; Cho et al., 2024; Wang et al., 2024d; Kuang et al., 2024; Wu et al., 2024; Bian et al., 2024b) have explored federated fine-tuning approaches. (Kuang et al., 2024) proposed federated fine-tuning with all parameters updated, while (Sun et al., 2022) introduced federated fine-tuning with PEFT using prefix-tuning. (Zhang et al., 2024) was the first study to apply LoRA in a federated context. (Bai et al., 2024; Cho et al., 2024) focused on federated fine-tuning where clients have different LoRA ranks, while (Wang et al., 2024d) proposed strategies for LoRA initialization to achieve better performance. Another studies (Bian et al., 2024b; Sun et al., 2024) addressed server aggregation bias in LoRA-based federated fine-tuning.
The most relevant works to ours focus on personalized federated fine-tuning with LoRA (Anonymous, 2025; Yang et al., 2024a; Qi et al., 2024). FedDPA (Yang et al., 2024a) introduced global and local LoRA components but sufferd from interference during global LoRA training and struggled to balance the two phases effectively. PF2LoRA (Anonymous, 2025) employed a bilevel framework to combine a shared adapter with client-specific adapters but faced interference issues during the aggregation of the shared adapter. FDLoRA (Qi et al., 2024) utilized personalized LoRA modules to initialize federated training, combining them with a global LoRA module via adaptive fusion; however, it relied on server-side datasets and was prone to interference during global LoRA aggregation, limiting its effectiveness in heterogeneous settings.
6 Conclusion
In this work, We proposed FedALT, a personalized federated LoRA fine-tuning framework that departs from the FedAvg paradigm by prioritizing local adaptation while integrating global knowledge via a Rest-of-the-World (RoTW) LoRA component. An adaptive mixer dynamically balances local and global information, mitigating cross-client interference. Experiments across NLP tasks demonstrate FedALT’s superiority in heterogeneous settings with minimal overhead. Future work will focus on a theoretical analysis of FedALT’s convergence.
Appendix A Datasets
In this paper, we primarily follow the setup of (Yang et al., 2024a) to construct federated datasets for heterogeneous client settings using the Flan datasets (Chung et al., 2024). The Flan collection consists of over 12 NLP tasks, each containing multiple datasets. These tasks, generated from diverse contextual factors, inherently experience various complex distribution shifts. We consider two federated dataset configurations. For the first configuration, we assume an extreme client heterogeneity setting, where 8 clients are assigned completely different NLP tasks (a total of 8 distinct tasks are considered). For the second configuration, we consider a relatively moderate setting, where some clients share the same NLP task but with data originating from different datasets. In this setting, 8 clients with 6 distinct NLP tasks are included. To simulate a realistic FL environment, where each client has limited access to large datasets, we employ a downsampling strategy, assigning each client 600 training samples and 300 testing samples. The specific tasks and datasets for each federated dataset configuration are detailed in Tables 5 and 6. Note that while there is some overlap in the datasets used for federated setting 1 and federated setting 2, the samples within each configuration are randomly selected, ensuring independence between the two settings.
The original NLP tasks in Flan exhibit diverse formats, making it challenging to standardize processing for LLMs and to select a consistent evaluation metric. Even within a single dataset (e.g., Sentiment140), samples can vary significantly in format, as illustrated in Tables 7 and 8. To facilitate uniform processing and evaluation, we adopt a simple prompt template following (Yang et al., 2024a), as shown in Table 9. ROUGE-1 is chosen as the evaluation metric to assess performance across all tasks.
Task Dataset Description
Text Classification ag_news_subset News classification into categories.
Natural Language Inference snli Determine relationships between sentence pairs.
Reading Comprehension openbookqa Open-book question answering.
Paraphrase Detection glue_mrpc Identify semantic equivalence in sentences.
Commonsense Reasoning story_cloze Choose the correct story ending.
Struct to Text common_gen Generate text from concepts.
Sentiment Analysis sentiment140 Sentiment analysis of tweets.
Coreference Resolution definite_pronoun_resolution Resolve ambiguous pronouns.
Table 5: Tasks and Datasets for Federated Dataset 1
Task Dataset Description
Struct to Text 1 common_gen Generate text from concepts.
Struct to Text 2 dart Generate structured text descriptions.
Coreference Resolution definite_pronoun_resolution Resolve ambiguous pronouns.
Natural Language Inference 1 snli Determine relationships between sentence pairs.
Natural Language Inference 2 qnli Answer questions based on context.
Paraphrase Detection glue_mrpc Identify semantic equivalence in sentences.
Text Classification ag_news_subset Classify news into categories.
Sentiment Analysis sst2 Analyze sentiment of text.
Table 6: Tasks and Datasets for Federated Dataset 2
Data Examples
Instruction: @celinieee The crafts thingy is already done. Describe the sentiment embodied by this tweet. OPTIONS: - negative - positive.
Output: negative
Dataset: sentiment140
Instruction: Write a positive tweet.
Output: @MandinaM Very well indeed thank you.. Yourself? Just about to start work.. 9 hours to go..
Dataset: sentiment140
Table 7: Examples of Data in Sentiment Analysis Task.
Data Examples
Instruction: What are the most important words in the following sentence: pig in a pen on a pig farm
Output: farm, pen, pig
Dataset: common_gen
Instruction: Generate a sentence about this data: Wildwood eatType coffee shop; Wildwood priceRange moderate; Wildwood customer rating 1 out of 5; Wildwood near Ranch
Output: The Wildwood coffee shop is an average-priced location that serves a light menu of food and alcohol. Located near the Ranch, customers give this location a 1 out of 5 stars.
Dataset: dart
Table 8: Examples of Data in Struct to Text Task.
Template
Prompt Input Instruction: {instruction} Response:
Table 9: Prompt Template.
Appendix B Training Implementations
We implement all methods using PyTorch and conduct all experiments on six NVIDIA A100 GPUs. For the federated methods, we set the local epochs to 5 and the global rounds to 20. For the Local Only baseline, the total number of training rounds is set to 50. The learning rate is set to $3\times 10^{-4}$, and the AdamW optimizer is utilized, following (Yang et al., 2024a).
LoRA modules are inserted into the $q$ and $v$ components of each self-attention layer in LLaMA2. The batch size is set to 8, and the rank of the trainable LoRA for each method is fixed at 8. Additionally, we set lora_alpha to 32 and lora_dropout to 0.05. Due to computational limitations, the pre-trained LLaMA2 model is loaded in 8-bit precision.
Appendix C Methods
Local Only. The Local Only method represents a setting where each client independently trains its own model without sharing any model updates with a server. This method serves as a baseline to assess the impact of federated learning methods on model performance.
FedIT (Zhang et al., 2024). FedIT integrates LoRA with the FedAvg framework, making it one of the earliest approaches to combine these techniques. However, under heterogeneous client data settings, FedIT suffers from harmful cross-client interference during the aggregation step, as highlighted in the motivational experiments. This interference often reduces its efficiency and, in some cases, results in worse performance than the Local Only method.
FFA-LoRA (Sun et al., 2024). FFA-LoRA aims to reduce communication costs by fixing the randomly initialized non-zero $\mathbf{A}$ matrices and fine-tuning only the zero-initialized $\mathbf{B}$ matrices. While this approach lowers communication overhead, it impairs the learning capacity of LoRA, as fixed $\mathbf{A}$ matrices limit the model’s ability to adapt to diverse client data, resulting in suboptimal performance.
FedSA (Guo et al., 2024). FedSA introduces an asymmetric approach to LoRA matrix updates in federated learning, where only the $\mathbf{A}$ matrices are shared with the server for aggregation, while the $\mathbf{B}$ matrices remain local to capture client-specific knowledge. However, even the client-specific $\mathbf{B}$ matrices may contain valuable knowledge that could benefit other clients. The failure to leverage this knowledge effectively results in reduced overall performance.
FedDPA (Yang et al., 2024a). FedDPA is a state-of-the-art personalized federated fine-tuning method that introduces two distinct LoRA components: global LoRA and local LoRA. The global LoRA is trained using FedAvg, while the local LoRA is independently fine-tuned on client-specific data. Despite its personalization capabilities, FedDPA suffers from interference during the global LoRA training stage. Additionally, its separation of local and global LoRA fine-tuning into distinct phases complicates the determination of optimal training durations for each phase.
PF2LoRA (Anonymous, 2025). PF2LoRA adopts a two-level low-rank adaptation framework, where a common adapter shared across all clients is combined with client-specific lightweight adapters for personalization. This bilevel optimization framework aims to balance shared knowledge and personalization. However, similar to FedDPA, the aggregation of the shared adapter during the server aggregation step can reintroduce harmful interference, limiting the method’s effectiveness in highly heterogeneous settings.
FDLoRA (Qi et al., 2024). FDLoRA follows a two-step process where personalized LoRA modules are first trained locally and then used to initialize federated training to obtain a global LoRA module. An adaptive fusion mechanism combines the two LoRA modules during fine-tuning. While this method improves personalization, it requires additional server-side datasets for effective global LoRA training. Moreover, the aggregation step during global LoRA training remains prone to interference, reducing its robustness in handling client heterogeneity.
Appendix D Impact of RoTW LoRA
To evaluate the impact of using the RoTW LoRA instead of the global average of all Local LoRAs (which includes each client’s Individual LoRA), we conduct an experiment. As shown in Figure 4, leveraging the RoTW LoRA significantly improves performance compared to using the global average. This result underscores the importance of isolating client-specific information from the aggregated global knowledge. By decoupling the RoTW LoRA from client updates, our method effectively mitigates cross-client interference, enabling better personalization.
Figure 4: Impact of RoTW LoRA. Results indicate that isolating client-specific information is more effective than directly using globally averaged information. See details in Appendix D.
Appendix E Discussions
E.1 Limitations of FedALT in Traditional FL
In FedALT, a separate RoTW LoRA is maintained alongside each client’s Individual LoRA component. The RoTW LoRA is not trained during each round but remains fixed, which slightly increases the forward/inference workload. However, we emphasize that FedALT is particularly well-suited for federated LoRA fine-tuning. Compared to the pre-trained model (e.g., LLaMA), the additional forward computation introduced by the RoTW LoRA is minimal and can be largely neglected.
However, applying the Individual and RoTW structure to traditional FL, where the full model is updated rather than a lightweight adapter, may not be practical. In such a setting, maintaining an additional RoTW model for inference would effectively double the forward computation workload, making it computationally inefficient. Therefore, our method is specifically designed for federated fine-tuning scenarios, where the overhead remains manageable and the benefits of interference isolation are maximized.
E.2 Comparison with FedDPA (Yang et al., 2024a)
FedDPA also introduces two LoRA components, referred to as global LoRA and local LoRA. By incorporating an additional local fine-tuning step excluded from the aggregation process, FedDPA aims to enhance personalization for individual clients. However, our proposed method, FedALT, differs from FedDPA in several key aspects: First, in both versions of FedDPA, global LoRA training is independent of local LoRA training. During global training, the local LoRA components are excluded, meaning that interference persists during the global training process, similar to FedIT. Furthermore, because the global LoRA training in FedDPA is independent of local fine-tuning, the local LoRA is never shared with other clients. This is in stark contrast to our method, where the RoTW LoRA captures useful training information from all other clients. Additionally, the dynamic $\alpha$ weight in FedALT allows each client to individually determine the contribution of its Individual LoRA and Rest-of-World LoRA on the per-input basis, enabling a more personalized and effective model. Second, FedDPA separates local and global LoRA fine-tuning into distinct steps, which introduces challenges in determining the duration of each step. If the local fine-tuning step is too short, the personalization effect is limited; if it is too long, the process approaches independent local fine-tuning, thereby negating the potential benefits of FL. Such additional hyperparameters can significantly affect performance but are often difficult to tune. In contrast, our proposed method eliminates the need to separate local and global fine-tuning phases and has no additional hyperparameters, simplifying the process and avoiding these timing-related challenges.

### v2

@@
-FedALT: Federated Fine-Tuning through Adaptive Local Training with Rest-of-the-World LoRA
+FedALT: Federated Fine-Tuning through Adaptive Local Training with Rest-of-World LoRA
 Abstract
@@
 In this work, we propose FedALT, a novel personalized federated LoRA fine-tuning algorithm that fundamentally departs from FedAvg.
-Instead of using an aggregated model to initialize local training, each client continues training its individual LoRA while incorporating shared knowledge through a separate Rest-of-the-World (RoTW) LoRA component.
-To effectively balance local adaptation and global information, FedALT introduces an adaptive mixer that dynamically learns input-specific weightings between the individual and RoTW LoRA components using the Mixture-of-Experts (MoE) principle.
+Instead of using an aggregated model to initialize local training, each client continues training its individual LoRA while incorporating shared knowledge through a separate Rest-of-World (RoW) LoRA component.
+To effectively balance local adaptation and global information, FedALT introduces an adaptive mixer that dynamically learns input-specific weightings between the individual and RoW LoRA components, drawing conceptual foundations from the Mixture-of-Experts (MoE) paradigm.
 Through extensive experiments on NLP benchmarks, we demonstrate that FedALT significantly outperforms state-of-the-art personalized federated LoRA fine-tuning methods, achieving superior local adaptation without sacrificing computational efficiency.
-Machine Learning, ICML
 1 Introduction
 Figure 1: Illustration of FedALT.
-Instead of directly aggregating local LoRA modules from each client using FedAvg, FedALT introduces a frozen Rest-of-World LoRA component to transmit shared global knowledge while preserving client-specific adaptations through Individual LoRA modules.
-The adaptive mixer dynamically combines the RoTW LoRA and Individual LoRA.
+Instead of directly aggregating local LoRA modules from each client using FedAvg, FedALT introduces a frozen RoW LoRA component to transmit shared global knowledge while preserving client-specific adaptations through Individual LoRA.
+The adaptive mixer dynamically combines the RoW LoRA and Individual LoRA.
 See details in section 3.
-Large language models (LLMs) (Kenton & Toutanova, 2019; Brown et al., 2020; Raffel et al., 2020; Touvron et al., 2023a; Zhou et al., 2024; Zeng et al., 2022; Touvron et al., 2023b) have demonstrated exceptional capabilities in language understanding and generation, enabling a wide range of applications.
-Adapting pretrained LLMs to specialized domains or further enhancing their performance requires fine-tuning with task-specific data (Minaee et al., 2024).
+Large language models (LLMs) [19, 4, 31, 37, 50, 48, 38] have demonstrated exceptional capabilities in language understanding and generation, enabling a wide range of applications.
+Adapting pretrained LLMs to specialized domains or further enhancing their performance requires fine-tuning with task-specific data [29].
 However, full fine-tuning, which involves updating all model parameters, is computationally expensive and often impractical for real-world deployment.
-To address this limitation, several parameter-efficient fine-tuning (PEFT) (Ding et al., 2023; Fu et al., 2023; Han et al., 2024; Liu et al., 2022; Li & Liang, 2021; Lester et al., 2021) methods have been proposed, with low-rank adaptation (LoRA) (Hu et al., 2021) being one of the most widely used.
+To address this limitation, several parameter-efficient fine-tuning (PEFT) [9, 11, 14, 25, 24, 23] methods have been proposed, with low-rank adaptation (LoRA) [16] being one of the most widely used.
 LoRA reduces the number of trainable parameters by integrating low-rank matrices into the model, significantly reducing computational costs.
-Fine-tuning LLMs requires substantial data to adapt to downstream tasks, yet this data is often distributed across multiple institutions where privacy regulations and security concerns prohibit direct sharing.
-Federated learning (FL) (McMahan et al., 2017) enables collaborative fine-tuning without exposing raw data, making it an essential framework for privacy-preserving adaptation.
-However, most existing federated LoRA fine-tuning approaches (Zhang et al., 2024; Sun et al., 2024; Wang et al., 2024d; Guo et al., 2024; Bian et al., 2024b) assume a single global model, overlooking the highly heterogeneous data distributions among clients.
-A one-size-fits-all model often results in suboptimal performance and fairness issues, especially in real-world scenarios where client data varies significantly in size, domain, and distribution.
-For example, in healthcare (Kim et al., 2024), hospitals may require fine-tuned models tailored to specific patient demographics or disease profiles.
-Without personalization, federated fine-tuning fails to capture such variations, leading to poor local adaptation and inefficient resource utilization.
-Despite the critical need for personalized solutions, personalized federated LoRA fine-tuning (Yang et al., 2024a; Anonymous, 2025; Qi et al., 2024) remains largely underexplored.
-Federated Averaging (FedAvg) (McMahan et al., 2017) is perhaps the most widely used FL framework, serving as the foundation for numerous follow-up works that introduce various optimizations.
+Fine-tuning LLMs necessitates substantial data volumes for adaptation to downstream tasks; however, this requisite data often resides across multiple institutions where privacy regulations and security protocols prohibit direct data sharing.
+Federated Learning (FL) [27] facilitates collaborative fine-tuning without raw data exposure, establishing itself as an essential paradigm for privacy-preserving adaptation.
+Nonetheless, existing federated LoRA fine-tuning methodologies [49, 34, 43, 12, 3] predominantly presuppose a singular global model, neglecting the inherent data heterogeneity among participating clients.
+Such a homogeneous modeling approach frequently yields suboptimal performance and introduces fairness disparities, particularly in real-world deployments where client data exhibits significant variance in volume, domain, and distributional characteristics.
+While limited research on personalized federated LoRA fine-tuning [15, 30] exists, these approaches primarily address label heterogeneity cases.
+However, in realistic LLM fine-tuning contexts, clients’ learning objectives naturally diverge due to diverse business imperatives [47].
+Personalized federated LoRA fine-tuning addressing more realistic task-heterogeneous scenarios [45] remains insufficiently investigated in the current literature.
+Federated Averaging (FedAvg) [27] is perhaps the most widely used FL framework, serving as the foundation for numerous follow-up works that introduce various optimizations.
 Its core principle is to aggregate locally trained client models into a global model, which then initializes the next round of local training.
-Most personalized FL algorithms (Yang et al., 2024a; Qi et al., 2024; Tan et al., 2022; Deng et al., 2020; Collins et al., 2021; Fallah et al., 2020), despite their implementation differences, build upon the FedAvg framework by incorporating personalization mechanisms—and personalized federated LoRA fine-tuning follows the same paradigm.
-For example, the recently proposed FedDPA (Yang et al., 2024a) algorithm personalizes LoRA by introducing both global and local LoRA components.
+Most personalized FL algorithms [45, 30, 35, 8, 7, 10], despite their implementation differences, build upon the FedAvg framework by incorporating personalization mechanisms—and personalized federated LoRA fine-tuning follows the same paradigm.
+For example, the recently proposed FedDPA [45] algorithm personalizes LoRA by introducing both global and local LoRA components.
 The global LoRA component is trained using FedAvg, while the local LoRA component is trained independently on each client’s private data.
 This local adaptation can occur either after the global LoRA stabilizes or alternatedly with global LoRA training, in which case the final output is computed as a weighted combination of both components using a pre-defined weighting factor.
-However, while personalization improves the average performance across clients, some clients may experience degraded performance, even falling behind simple local fine-tuning—that is, not participating in FL at all.
-This issue, also observed in our motivational study in Section 2.3, raises a critical question: Is FedAvg truly the most effective framework for developing personalized federated LoRA fine-tuning methods in heterogeneous settings?
-In this paper, we propose a novel personalized federated LoRA fine-tuning algorithm that departs fundamentally from the FedAvg paradigm.
-Our method is motivated by identifying two key limitations of FedAvg-based approaches, such as FedDPA, in handling FL challenges under heterogeneous client settings.
-First, FedAvg-based approaches suffer from harmful cross-client interference due to their reliance on a global model aggregated from client models.
-This aggregation step, intended to capture shared knowledge, often disrupts individual client adaptations when data distributions are substantially different.
-As a result, any improvements achieved during local fine-tuning can be negated by aggregation, making it ineffective for clients whose distributions deviate significantly from the global model.
-For such clients, initializing the next training round with the aggregated model may offer little to no benefit—or even degrade performance.
-Second, FedAvg-based approaches lack an effective mechanism to balance local and global information in personalization.
-For example, FedDPA combines global and local LoRA components using a pre-defined weighting factor during training.
-However, improper weighting can significantly harm personalized fine-tuning, particularly when the global model is assigned excessive weight.
-This issue is especially problematic for clients where the global model is poorly suited, leading to suboptimal performance despite the presence of a local adaptation mechanism.
-Overall, the FedAvg paradigm structures the personalization mechanisms around the aggregated model, treating local training as a supplement to the global model.
-In contrast, our approach emphasizes the role of local training in heterogeneous settings, focusing on personalization by incorporating useful global information directly into local training.
-Specifically, we eliminate the key FedAvg step of using the aggregated model to initialize each local training round.
-Instead, clients continue training on their own local models from the previous round.
-Information from other clients’ models is aggregated into a “rest-of-the-world” (RoTW) LoRA component, which supports local training but remains frozen during each round.
-Furthermore, to effectively leverage the RoTW component, we introduce an “adaptive mixer” that dynamically learns the optimal weighting between the individual LoRA and RoTW LoRA components using the Mixture-of-Experts (MoE) principle (Jordan & Jacobs, 1994).
-This adaptive weighting enables the model to selectively integrate relevant global knowledge while maintaining local model adaptation.
-We name our approach FedALT, short for Federated Fine-tuning through Adaptive Local Training with Rest-of-the-World LoRA.
-We highlight that FedALT is particularly well-suited for federated LoRA fine-tuning due to its low computational complexity and minimal memory requirements.
-Although a separate RoTW LoRA is maintained alongside each client’s individual LoRA component, the RoTW LoRA is not trained during each round.
-Furthermore, LoRA modules are significantly smaller than the full model, meaning the added memory overhead is negligible.
-Through comprehensive experiments on heterogeneous FL benchmarks across various NLP tasks, we demonstrate that FedALT outperforms state-of-the-art approaches by a significant margin.
+However, our motivational study (see section 2.3) reveals that FedAvg-based approaches may degrade client-specific performance due to harmful cross-client interference.
+We identify two fundamental limitations of FedAvg-based methods like FedDPA in heterogeneous FL settings.
+First, they suffer from harmful cross-client interference due to reliance on global aggregation, which can negate individual client improvements achieved during local fine-tuning.
+Second, these methods lack an effective mechanism for balancing global and local information, relying on fixed or predefined weighting factors, which can be suboptimal for clients poorly represented by the global model.
+To address these critical limitations, we propose a novel personalized federated LoRA fine-tuning approach called FedALT (short for Federated Fine-tuning through Adaptive Local Training with Rest-of-World LoRA.).
+Distinctly different from FedAvg-based methods, FedALT does not aggregate local models to initialize training rounds.
+Instead, each client continues training on their own previously trained local models, supplemented by a frozen Rest-of-World (RoW) LoRA component.
+The RoW LoRA captures global knowledge without updating during local training rounds, thus avoiding interference with client-specific adaptations.
+To leverage RoW LoRA effectively, we introduce an adaptive mixer, inspired by Mixture-of-Experts (MoE) [18], which dynamically and optimally combines global and local information for each client’s data sample.
+Our contributions are as follows:
+  • We highlight that personalized federated fine-tuning for clients with heterogeneous tasks has been insufficiently explored and identify critical limitations in existing methods that build upon the FedAvg paradigm.
+  • We propose FedALT, a novel personalization framework featuring a frozen RoW LoRA and an adaptive mixer, effectively addressing these limitations by dynamically balancing global and local model adaptations.
+  • We perform comprehensive experiments on two widely-used LLMs (Bloom [21] and Llama 2 [38]) using the Flan benchmark [6], demonstrating that FedALT significantly outperforms existing federated LoRA fine-tuning methods.
 2 Preliminaries and Motivations
@@
 The objective of the training process can be formulated as:
-$\min_{\mathcal{W}}\frac{1}{K}\sum_{k=1}^{K}L(X_{k},Y_{k},\mathbf{W}_{k}),$ (2)
-where $L(\cdot)$ represents the loss function, and $\mathcal{W}$ denotes the set of personalized fine-tuned models $\{\mathbf{W}_{k}\}_{k=1}^{K}$.
+$\min_{\mathcal{W}}\frac{1}{K}\sum_{k=1}^{K}L_{k}(X_{k},Y_{k},\mathbf{W}_{k}),$ (2)
+where $L_{k}(\cdot)$ represents the loss function for client $k$, and $\mathcal{W}$ denotes the set of personalized fine-tuned models $\{\mathbf{W}_{k}\}_{k=1}^{K}$.
 2.3 Motivational Study
-Federated fine-tuning operates in highly heterogeneous environments where client datasets can differ significantly.
+Federated fine-tuning operates in highly heterogeneous environments where client tasks can differ significantly.
 To understand the effects of such heterogeneity, we conducted a motivational study that highlights both the benefits and challenges of applying FL in these settings.
-Figure 2: Motivational study results shows that while FedIT underperforms Local Only on some clients due to harmful cross-client interference.
-FedIT-split, using multiple smaller LoRAs, fails to mitigate this interference, highlighting the limitations of the FedAvg paradigm.
-See details in Section 2.3.
 Fact 1: FL Can Transfer Useful Knowledge but Also Introduces Harmful Interference.
-In FL settings, the data collected by each client can be highly diverse (Huang et al., 2022; Mendieta et al., 2022; Wang et al., 2024b).
-We conducted a motivational experiment using the Flan dataset (Chung et al., 2024), selecting 8 tasks and creating 8 clients, each specializing in a single task.
+In FL settings, the data collected by each client can be highly diverse [17, 28, 40].
+We conducted a motivational experiment using the Flan dataset [6], selecting 8 tasks and creating 8 clients, each specializing in a single task.
 To demonstrate potential interference in federated fine-tuning, we compared two baseline methods.
-The first method, FedIT (Zhang et al., 2024), combines FedAvg with LoRA fine-tuning.
+The first method, FedIT [49], combines FedAvg with LoRA fine-tuning.
 The second method is Local Only LoRA fine-tuning, where each client fine-tunes its LoRA locally without any communication or information sharing between clients.
-The results, shown in Figure 2, support our first observation.
+Figure 2: Motivational study results. See details in section 2.3.
+The results, shown in fig.
+2, support our first observation.
 On one hand, federated fine-tuning using FedIT achieves better performance than local fine-tuning on tasks Coreference Resolution and Paraphrase Detection, confirming that FL can enhance learning by enabling clients to benefit from shared knowledge.
@@
 Having established the existence of harmful interference, we next explored potential mitigation strategies.
-In centralized LoRA fine-tuning, HydraLoRA (Tian et al., 2024) found that splitting a single large LoRA into multiple smaller LoRAs reduces cross-domain interference.
+In centralized LoRA fine-tuning, HydraLoRA [36] found that splitting a single large LoRA into multiple smaller LoRAs reduces cross-domain interference.
 However, FL introduces additional challenges, as data domains are distributed across clients.
@@
 FedIT-split).
-However, as shown in Figure 2, its performance remained comparable to FedIT (FedAvg with a single large LoRA), indicating that interference persisted.
+However, as shown in fig.
+2, its performance remained comparable to FedIT (FedAvg with a single large LoRA), indicating that interference persisted.
 This result is expected: in centralized settings, interference occurs within a shared dataset and can be mitigated by partitioning LoRAs.
@@
 3 Proposed Method: FedALT
-Our motivational studies reveal that harmful interference in federated learning arises from server aggregation and the use of the aggregated model to initialize subsequent training rounds.
+Our motivational studies reveal that harmful interference in FL arises from server aggregation and the use of the aggregated model to initialize subsequent training rounds.
 In the FedAvg paradigm, a global LoRA serves as a shared initialization point, allowing clients to fine-tune locally on their respective datasets.
 After fine-tuning, the updated LoRAs are averaged on the server and redistributed to clients as the new initialization.
-However, when client data distributions differ significantly, aggregation disrupts individual adaptations, often negating improvements from local training and hindering personalization.
+However, when client data differ significantly, aggregation disrupts individual adaptations, often negating improvements from local training and hindering personalization.
 Can cross-client interference be mitigated?
@@
 Our method introduces two key innovations: 1.
-Each client maintains two LoRA modules: an Individual LoRA, which captures locally learned information and is updated exclusively during the client’s fine-tuning, and a Rest-of-the-World (RoTW) LoRA, which aggregates information from all other clients.
-Crucially, the RoTW LoRA remains fixed during local fine-tuning, ensuring that aggregation does not interfere with a client’s adaptation.
+Each client maintains two LoRA modules: an Individual LoRA, which captures locally learned information and is updated exclusively during the client’s fine-tuning, and a Rest-of-World (RoW) LoRA, which aggregates information from all other clients.
+Crucially, the RoW LoRA remains fixed during local fine-tuning, ensuring that aggregation does not interfere with a client’s adaptation.
 2.
-An adaptive mixer dynamically learns the optimal input-specific weighting between the Individual LoRA and RoTW LoRA components, leveraging the Mixture-of-Experts (MoE) principle to balance personalization and global knowledge integration.
+An adaptive mixer dynamically learns the optimal input-specific weighting between the Individual LoRA and RoW LoRA components to balance personalization and global knowledge integration.
 In the following of this section, we detail the design and functionality of these components.
-3.1 Individual and Rest-of-the-World (RoTW) LoRA
-Each client $k$ maintains two types of LoRAs: an Individual LoRA, denoted by $\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$, and a Rest-of-the-World (RoTW) LoRA, denoted by $\mathbf{A}^{R}_{k}/\mathbf{B}^{R}_{k}$.
+Additionally, a theoretical convergence analysis of the proposed method is provided in the appendix F.
+3.1 Individual and Rest-of-World (RoW) LoRA
+Each client $k$ maintains two types of LoRAs: an Individual LoRA, denoted by $\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$, and a Rest-of-World (RoW) LoRA, denoted by $\mathbf{A}^{R}_{k}/\mathbf{B}^{R}_{k}$.
 The Individual LoRA captures information obtained locally by the client and is isolated from the information of other clients.
-In contrast, the RoTW LoRA represents the average of the Individual LoRAs from all other clients and remains frozen during the client’s local fine-tuning.
+In contrast, the RoW LoRA represents the average of the Individual LoRAs from all other clients and remains frozen during the client’s local fine-tuning.
 Specifically, at the end of each round, client $k$ updates its Individual LoRA $\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$ through local training.
-The RoTW LoRA for client $k$ is then computed as:
+The RoW LoRA for client $k$ is then computed as:
 $\mathbf{A}^{R}_{k}=\frac{1}{K-1}\sum_{m\neq k}\mathbf{A}^{L}_{m},\quad\mathbf{B}^{R}_{k}=\frac{1}{K-1}\sum_{m\neq k}\mathbf{B}^{L}_{m},$ (3)
-After clients upload their local LoRAs to the server, the server computes the RoTW LoRA for each client individually and distributes it back to them.
+After clients upload their local LoRAs to the server, the server computes the RoW LoRA for each client individually and distributes it back to them.
 Alternatively, the server can calculate the global average of all clients’ individual LoRAs and distribute this average back to the clients.
-Each client then determines its RoTW LoRA by subtracting its individual LoRA from the global average LoRA.
-Now, consider the scenario where a weight $\frac{1}{K}$ is assigned to the local LoRA and $1-\frac{1}{K}$ to the RoTW LoRA when combining their outputs.
+Each client then determines its RoW LoRA by subtracting its individual LoRA from the global average LoRA.
+Now, consider the scenario where a weight $\frac{1}{K}$ is assigned to the local LoRA and $1-\frac{1}{K}$ to the RoW LoRA when combining their outputs.
 In this case, the combined LoRA is mathematically equivalent to the averaged LoRA.
-This equivalence arises naturally from the definitions of the RoTW and averaged LoRAs.
+This equivalence arises naturally from the definitions of the RoW and averaged LoRAs.
 If these two LoRAs are trained simultaneously on local datasets, the approach essentially reduces to the standard FedAvg (or FedIT) algorithm, as all clients would end up with the same combined LoRA.
-However, it is crucial to emphasize that the RoTW LoRA is not further updated on clients using their local datasets.
+However, it is crucial to emphasize that the RoW LoRA is not further updated on clients using their local datasets.
 Instead, it is only updated through averaging at the end of each learning round.
-Training the RoTW LoRA on local datasets is unnecessary and potentially counterproductive, as subsequent averaging could cancel out any improvements.
-Furthermore, skipping local training for the RoTW LoRA reduces the computational workload on clients by half.
+Training the RoW LoRA on local datasets is unnecessary and potentially counterproductive, as subsequent averaging could cancel out any improvements.
+Furthermore, skipping local training for the RoW LoRA reduces the computational workload on clients by half.
 3.2 Dynamic Mixture-of-Experts Mechanism
-A natural question arises: why not simply add the RoTW LoRA directly to the pre-trained model, i.e., $\mathbf{W}_{0}\leftarrow\mathbf{W}_{0}+\mathbf{B}^{R}_{k}\mathbf{A}^{R}_{k}$, since the RoTW LoRA remains fixed during local training?
+A natural question arises: why not simply add the RoW LoRA directly to the pre-trained model, i.e., $\mathbf{W}_{0}\leftarrow\mathbf{W}_{0}+\mathbf{B}^{R}_{k}\mathbf{A}^{R}_{k}$, since the RoW LoRA remains fixed during local training?
 This approach would allow each client to maintain only a single Individual LoRA, saving memory space.
@@
   1. 1.
-    Potential Contamination of the Pre-trained Model: If the RoTW LoRA underperforms, adding it directly to the pre-trained model risks contaminating it, making further corrections difficult.
+    Potential Contamination of the Pre-trained Model: If the RoW LoRA underperforms, adding it directly to the pre-trained model risks contaminating it, making further corrections difficult.
   2. 2.
-    Loss of Flexibility: Directly incorporating the RoTW LoRA enforces a fixed weight, preventing dynamic adjustment between the Individual LoRA and shared information for different inputs.
+    Loss of Flexibility: Directly incorporating the RoW LoRA enforces a fixed weight, preventing dynamic adjustment between the Individual LoRA and shared information for different inputs.
 This lack of flexibility is problematic because different inputs may benefit variably from the local model versus the globally averaged model.
 A “one-size-fits-all” approach often fails to perform optimally across diverse inputs.
-To address these challenges, we propose a dynamic weighting mechanism for combining the Individual LoRA and the RoTW LoRA, enabling input-specific flexibility.
+To address these challenges, we propose a dynamic weighting mechanism for combining the Individual LoRA and the RoW LoRA, enabling input-specific flexibility.
 Specifically, we introduce a mixer to dynamically adjust the contributions of the two LoRAs for each input.
@@
 $y=\mathbf{W}_{0}x+\alpha_{k}(x)\mathbf{B}^{L}_{k}\mathbf{A}^{L}_{k}x+\big{(}1-\alpha_{k}(x)\big{)}\mathbf{B}^{R}_{k}\mathbf{A}^{R}_{k}x,$
-where $\alpha_{k}(x)$ determines for client $k$ the contribution weight of the Individual LoRA, and $1-\alpha_{k}(x)$ determines the contribution of the RoTW LoRA.
+where $\alpha_{k}(x)$ determines for client $k$ the contribution weight of the Individual LoRA, and $1-\alpha_{k}(x)$ determines the contribution of the RoW LoRA.
 Given the complexity of modern LLMs, where LoRA modules are inserted at multiple layers, learning $\alpha(x)$ is non-trivial.
-Inspired by (Jordan & Jacobs, 1994), we adopt a Mixture-of-Experts (MoE) approach to compute $\alpha(x)$.
+Inspired by [18], we adopt a Mixture-of-Experts (MoE) approach to compute $\alpha(x)$.
 Specifically, we introduce a mixer $\mathbf{G}_{k}\in\mathbb{R}^{2\times d}$ which is a dense layer with trainable weights (a transformation matrix), followed by a softmax function, to learn the weighting:
@@
 3.3 Algorithm Workflow
-By combining the two LoRA modules (Individual LoRA and RoTW LoRA) with dynamic weights, our proposed method achieves a balance between leveraging useful information from other clients and isolating harmful interference.
+By combining the two LoRA modules (Individual LoRA and RoW LoRA) with dynamic weights, our proposed method achieves a balance between leveraging useful information from other clients and isolating harmful interference.
 To better illustrate the FedALT training process, we provide an overview of the server and client operations during fine-tuning at round $t$.
@@
 After local fine-tuning in round $t$, the server receives each client $k$’s Individual LoRA modules ($\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$).
-Upon receiving all clients’ Individual LoRA modules, the server updates and computes the RoTW LoRA modules for each client $k$ based on Equation 3.
-After calculating the RoTW LoRA modules for client $k$, the server broadcasts these updated RoTW LoRA modules to the respective client.
+Upon receiving all clients’ Individual LoRA modules, the server updates and computes the RoW LoRA modules for each client $k$ based on eq.
+3.
+After calculating the RoW LoRA modules for client $k$, the server broadcasts these updated RoW LoRA to the respective client.
 Client Side.
-Each client $k$ replaces its current RoTW LoRA modules with the new RoTW LoRA modules ($\mathbf{A}^{R}_{k}/\mathbf{B}^{R}_{k}$) received from the server.
+Each client $k$ replaces its current RoW LoRA modules with the new RoW LoRA modules ($\mathbf{A}^{R}_{k}/\mathbf{B}^{R}_{k}$) received from the server.
 It then begins local fine-tuning for round $t+1$ using its local dataset.
-During the local fine-tuning process, the client updates the parameters in its Individual LoRA modules ($\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$) and the mixer $\mathbf{G}_{k}$, while the RoTW LoRA modules ($\mathbf{A}^{R}_{k}/\mathbf{B}^{R}_{k}$) and the pre-trained model ($\mathbf{W}_{0}$) remain fixed.
+During the local fine-tuning process, the client updates the parameters in its Individual LoRA modules ($\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$) and the mixer $\mathbf{G}_{k}$, while the RoW LoRA modules ($\mathbf{A}^{R}_{k}/\mathbf{B}^{R}_{k}$) and the pre-trained model ($\mathbf{W}_{0}$) remain fixed.
 Once the local fine-tuning is complete, the updated Individual LoRA is uploaded to the server, while $\mathbf{G}_{k}$ stays local to the client.
-3.4 Discussion: Why a Single RoTW LoRA?
+The theoretical analysis of the convergence of the proposed FedALT is presented in appendix F.
+3.4 Discussion: Why a Single RoW LoRA?
 FedALT prioritizes local training while selectively incorporating useful information from other clients via a dynamic weighting mechanism.
@@
 $y=\mathbf{W}_{0}x+\alpha_{k,k}(x)\mathbf{B}^{L}_{k}\mathbf{A}^{L}_{k}x+\sum_{j\neq k}\alpha_{k,j}(x)\mathbf{B}^{L}_{j}\mathbf{A}^{L}_{j}x$
-where $\sum_{j}\alpha_{k,j}(x)=1$. This would allow client $k$ to capture the distinct influence of each other client’s LoRA.
+where $\sum_{j}\alpha_{k,j}(x)=1$.
+This would allow client $k$ to capture the distinct influence of each other client’s LoRA.
 However, this approach presents significant drawbacks.
@@
 Second, training the mixer becomes increasingly complex with more LoRA components, requiring additional parameters and longer convergence times.
-To address these challenges, we adopt a single RoTW LoRA, which aggregates knowledge from other clients into a single component.
+To address these challenges, we adopt a single RoW LoRA, which aggregates knowledge from other clients into a single component.
 While this reduces computational and communication costs, it may limit personalization.
-Future work could explore clustering-based approaches to maintain a small number of RoTW LoRA components, capturing finer-grained client relationships while remaining scalable.
+Future work could explore clustering-based approaches to maintain a small number of RoW LoRA components, capturing finer-grained client relationships while remaining scalable.
 4 Experiments
-In this section, we evaluate and compare the performance of our proposed method, FedALT, with other baseline methods across two federated learning dataset settings.
-All experiments are conducted on four NVIDIA A100 GPUs.
-Detailed training implementations are provided in the Appendix B.
-The code and datasets used for the experiments will be made publicly available upon acceptance of the paper.
+In this section, we evaluate and compare the performance of our proposed method, FedALT, with other baseline methods. Detailed training implementations are provided in the appendix B.
 Datasets.
-We construct our training datasets from Flan (Chung et al., 2024), a comprehensive collection of over 60 NLP datasets covering more than 12 diverse NLP tasks.
-For our experiments, we consider two federated dataset settings.
-In the first setting, we simulate extreme client heterogeneity by assuming 8 clients, each having a completely different NLP task dataset.
-In the second setting, we assume 8 clients with datasets spanning a total of 6 NLP tasks.
-Unlike the first setting, some clients in the second setting share the same NLP task, even though their datasets come from different sources.
-Detailed descriptions of the datasets, including task distributions and client assignments, are provided in the Appendix A.
+Our training datasets are derived from Flan [6], a comprehensive benchmark comprising over 60 NLP datasets across more than 12 diverse task types.
+To simulate client heterogeneity in the main experiments, we assume 8 clients, each assigned a distinct NLP task dataset.
+Sensitivity analysis with varying numbers of clients and degrees of task heterogeneity are presented in section 4.3.
+Detailed dataset descriptions, including task distributions and client assignments, are provided in appendix A.
 Pre-trained Model and Baselines.
-We utilize LLaMA2-7B (Touvron et al., 2023b) as the pre-trained model for all baseline and proposed fine-tuning methods.
-To evaluate the performance of our proposed method, we compare it with the state-of-the-art methods in federated fine-tuning with LoRA, including FedIT (Zhang et al., 2024), FFA-LoRA (Sun et al., 2024), FedSA (Guo et al., 2024).
-Additionally, we compare it with state-of-the-art methods focused on personalized federated fine-tuning with LoRA, such as FedDPA (Yang et al., 2024a), PF2LoRA (Anonymous, 2025), and FDLoRA (Qi et al., 2024).
+We utilize LLaMA2-7B [38] and Bloom-560M [21] as the pre-trained models for all baseline and proposed fine-tuning methods.
+To evaluate the performance of our proposed method, we compare it with the state-of-the-art methods in federated fine-tuning with LoRA, including FedIT [49], FFA-LoRA [34].
+Additionally, we compare it with state-of-the-art methods focused on personalized federated fine-tuning with LoRA, such as FedDPA [45], PF2LoRA [15], FedSA [12], and FDLoRA [30].
 We also include the Local Only baseline, where clients fine-tune their models independently without federated collaboration.
 Detailed descriptions of these baseline methods are provided in the appendix C.
-Methods Dataset 1
+Table 1: Performance comparison with baseline methods with LLaMA2-7B. The average column represents the overall performance across these tasks. See details in section 4.1.
+Methods LLaMA2-7B
 Commonsense Coreference Natural Language Paraphrase Reading Sentiment Struct to Text Average
 Reasoning Resolution Inference Detection Comprehension Analysis Text Classification
-Local Only 73.83 74.62 58.73 71.00 55.82 46.56 55.14 67.18 62.81
-FedIT 72.82 77.14 59.53 72.33 53.35 46.76 49.23 66.39 62.22
-FFA-LoRA 66.96 68.48 50.41 69.79 49.69 44.44 43.29 61.15 55.54
+Local 73.83 74.62 58.73 71.00 55.82 46.56 55.14 67.18 62.86
+FedIT 72.82 77.14 59.53 72.33 53.35 46.76 49.23 66.39 62.19
+FFA-LoRA 66.96 68.48 50.41 69.79 49.69 44.44 43.29 61.15 56.78
 FedSA 72.68 78.24 64.46 76.33 54.32 42.71 53.57 65.47 63.47
@@
 FedALT 76.12 83.04 67.73 77.80 59.41 51.57 53.12 71.60 67.55
-Table 1: Performance comparison with baseline methods in the first federated dataset setting. The average column represents the overall performance across these tasks. See details in Section 4.1.
-Methods Dataset 2
-Struct to Struct to Coreference Natural Language Natural Language Paraphrase Text Sentiment Average
-Text 1 Text 2 Resolution Inference 1 Inference 2 Detection Classification Analysis
+Table 2: Performance comparison across various NLP tasks with Bloom-560M. The average column represents the overall performance across these tasks. See details in section 4.1.
+Methods Bloom-560M
+Commonsense Coreference Natural Language Paraphrase Reading Sentiment Struct to Text Average
+Reasoning Resolution Inference Detection Comprehension Analysis Text Classification
+Local 52.13 39.20 37.55 62.67 29.45 42.91 41.33 49.99 44.40
+FedIT 50.24 40.08 38.51 67.00 27.98 40.34 38.95 51.52 44.33
+FFA-LoRA 46.07 35.05 33.57 66.12 25.52 37.13 35.43 44.63 40.44
+FedSA 55.79 41.65 41.12 69.33 27.90 43.72 41.27 55.03 46.98
+FedDPA 55.43 41.88 41.80 69.80 29.74 41.89 39.54 55.85 46.99
+PF2LoRA 54.58 40.16 37.40 70.08 25.78 42.36 37.90 58.67 45.87
+FDLoRA 53.65 40.22 37.27 68.89 29.95 42.89 35.39 56.05 45.54
+FedALT 56.39 41.27 44.45 70.67 30.62 43.04 39.24 59.03 48.09
+4.1 Performance Comparisons
+We compare the performance of FedALT, with other baseline methods based on each client’s performance on its specific test dataset, aligned with its training dataset.
+As shown in table 1, which reports the results on LLaMA2-7B pretrained models, our proposed method achieves the best overall performance compared to all baselines, with the average score exceeding that of the best-performing baseline, FDLoRA, by 2.38.
+FedALT consistently delivers superior performance across most tasks, outperforming the baselines by at least 1.61 on the Natural Language Inference task and 2.02 on the Reading Comprehension task.
+We find that on the Struct-to-Text, the Local Only method achieves the highest performance.
+This is likely due to the unique nature of this task, making it difficult to benefit from shared global knowledge.
+Furthermore, we observe that FedIT underperforms compared to Local Only on four client tasks, indicating the presence of harmful interference during aggregation.
+Personalized federated LoRA fine-tuning methods also fall short compared to FedALT, likely because they fail to effectively isolate harmful interference during the server aggregation step.
+Despite their designs for local/global LoRA interaction, these methods update LoRA modules using aggregated global information, which may carry conflicting information from other clients.
+To further evaluate the robustness of our proposed FedALT, we conduct experiments using the Bloom-560M model.
+As shown in table 2, FedALT consistently outperforms all baseline methods across the NLP tasks.
+These empirical results validate that conventional FedAvg-based approaches suffer from harmful cross-client parameter interference during aggregation.
+In contrast, FedALT’s adaptive strategy, which balances local specialization and global knowledge, effectively alleviates this issue.
+The consistent performance gains observed with both LLaMA2-7B and Bloom-560M confirm the generalizability of our approach across diverse model families and parameter scales.
+4.2 Ablation Study
+Table 3: Impact of decoupling LoRA training. See Sec. 4.2.
+Methods Same Trainable Same Inference Avg. Perf. RoW-Update (rank = 4) $\checkmark$ $\times$ 63.59 RoW-Update (rank = 8) $\times$ $\checkmark$ 65.75 FedALT 67.55
+Impact of Decoupling LoRA Training.
+In our proposed method, only the Individual LoRA modules are updated locally during client fine-tuning, while the RoW LoRA remains frozen and is updated exclusively through server-side aggregation.
+This design mitigates harmful cross-client interference by preventing local updates to the shared global component.
+To evaluate the effectiveness of this decoupled training strategy, we compare FedALT with two alternative methods.
+The first baseline, RoW-Update (rank = 4), fine-tunes both the Individual and RoW LoRA modules locally, each with a rank of 4 to maintain the same total number of trainable parameters.
+The second baseline, RoW-Update (rank = 8), also fine-tunes both LoRA modules locally, but assigns a rank of 8 to each, resulting in a higher overall parameter count.
+For fair comparison, both baselines use the dynamic mixer.
+The key distinction lies in whether the RoW LoRA is updated locally.
+As shown in table 3, while the second method outperforms the first due to its increased capacity, both alternatives underperform relative to FedALT.
+Notably, even with more trainable parameters, the second method fails to match FedALT’s performance, highlighting the benefits of decoupling local updates from the global component.
+Impact of RoW LoRA To evaluate the impact of using the RoW LoRA instead of the global average of all Local LoRAs (which includes each client’s Individual LoRA), we conduct an experiment.
+As shown in fig.
+3(a), leveraging the RoW LoRA significantly improves performance compared to using the global average.
+This result underscores the importance of isolating client-specific information from the aggregated global knowledge.
+Impact of Dynamic Weighting via MOE.
+To enable input-specific flexibility, our proposed method adopts a MOE mechanism to dynamically adjust the contributions of the Individual LoRA and RoW LoRA for each data sample.
+We conduct experiments to evaluate the efficiency of this dynamic weighting by comparing it to two fixed weighting strategies.
+In the first fixed strategy, we set the contribution weight to a fixed value of $\alpha=0.5$, meaning that the Individual LoRA and RoW LoRA contribute equally.
+In the second fixed strategy, we set the contribution weight to $\alpha=1/K$, where $K$ is the number of clients.
+This configuration effectively reduces the method to using the averaged LoRA, relying solely on each client’s last-round LoRA modules for the next round’s initialization.
+The results in fig.
+3(b) validate that our proposed dynamic weighting mechanism outperforms both fixed strategies, demonstrating its ability to enhance performance.
+Additionally, our dynamic weighting is implemented through the personal $\mathbf{G}_{k}$ mixer, which is maintained locally by each client and is not aggregated during the training process.
+To further investigate this design choice, we conducted experiments where $\mathbf{G}_{k}$ is averaged across clients.
+The results confirm that retaining a personalized $\mathbf{G}_{k}$ for each client yields optimal performance, emphasizing the importance of input-specific and client-specific adjustments.
+(a) Impact of RoW LoRA. Isolating client-specific information is more effective than using globally averaged updates. See section 4.2.
+(b) Impact of dynamic weighting via MOE. The personalized mixer outperforms fixed or shared weighting. See section 4.2.
+Figure 3: Ablation studies of FedALT.
+4.3 Sensitivity Analysis
+Effect of Data Heterogeneity.
+To further evaluate the robustness of our method under varying degrees of client heterogeneity, we design experiments across three settings.
+Our main experimental configuration, where each client is assigned a unique task, is denoted as the High Heterogeneous setting.
+We also include two alternative configurations.
+In the Low Heterogeneous setting, 8 clients are divided between only two tasks—Sentiment Analysis and Natural Language Inference—with clients sharing a task drawing data from the same source, resulting in minimal heterogeneity.
+The Mild Heterogeneous setting involves 8 clients assigned datasets spanning a total of 6 distinct NLP tasks.
+As shown in table 4, FedALT consistently outperforms all baseline methods across all levels of heterogeneity, using both Bloom and LLaMA2 as pre-trained models.
+These results highlight the robustness and adaptability of FedALT in handling diverse and heterogeneous data distributions.
+Due to space limitations, we report only the average results here; detailed per-client performance can be found in the appendix E.
+Effect of Number of Clients.
+In our main experiments, we demonstrated the effectiveness of the proposed method using 8 clients.
+Here we extend our evaluation by conducting additional experiments with 16 and 24 clients, using BLOOM as the foundation model.
+For the 16-client setting, we distributed each original dataset from the 8-client configuration between two clients, resulting in each client having approximately half the data samples compared to the 8-client scenario, while maintaining the same 8 distinct task types across the federation.
+The results in fig.
+4(a) show that FedALT consistently maintains its performance advantage even as the number of clients increases, outperforming the strongest baseline by 1.90 at least.
+Note that FedALT remains efficient due to its architectural design: each client maintains only its Individual LoRA and a single RoW LoRA, rather than storing separate components for every other client.
+Additionally, the server-side computation of the RoW LoRA remains lightweight, as it involves only simple averaging operations.
+Table 4: Performance under varying heterogeneity levels. Evaluation on Bloom-560M and LLaMA2-7B models across three levels of task heterogeneity. See section 4.3.
+Methods Bloom-560M LLaMA2-7B Low Mild High Low Mild High Local 48.65 50.21 44.40 60.59 66.67 62.86 FedIT 52.15 49.03 44.33 62.24 64.53 62.19 FFA-LoRA 51.33 47.32 40.44 61.07 62.15 56.78 FedSA 53.03 52.55 46.98 62.87 67.44 63.47 FedDPA 53.12 52.06 46.99 62.06 67.49 64.64 PF2LoRA 51.10 51.97 45.87 63.56 67.60 64.44 FDLoRA 52.51 51.42 45.54 62.42 66.56 65.17 FedALT 54.48 54.89 48.09 64.66 70.48 67.55
+Effect of LoRA Rank.
+We evaluate the impact of different LoRA ranks $r\in\{4,6,8\}$ using BLOOM as the pre-trained model, keeping all other settings consistent with the main experiment.
+As shown in fig.
+4(b), the proposed method consistently outperforms baseline methods across all rank values, demonstrating its adaptability and robustness.
+Effect of Local Epochs.
+We also investigate the effect of varying local training epochs (5 and 10), while keeping all other configurations unchanged.
+As presented in fig.
+4(c), our method outperforms all baselines under both settings, further highlighting its robustness to different local training epochs.
+(a) Effect of number of clients. See section 4.3.
+(b) Effect of LoRA rank. See section 4.3.
+(c) Effect of local epochs. See section 4.3.
+Figure 4: Sensitivity Analyses of FedALT under different training configurations.
+5 Related Work
+FedAvg [27] addresses privacy and communication efficiency in federated learning by aggregating locally trained model parameters to construct a shared global model.
+With the growing interest in LLMs, FL has emerged as a promising framework for fine-tuning LLMs while preserving data privacy.
+Recent works have explored federated fine-tuning of LLMs by either updating the entire model [20] or applying PEFT techniques such as prefix-tuning [33] and LoRA [49].
+Subsequent studies [1, 5, 43, 3, 34] have tackled key challenges in this domain, including heterogeneous LoRA configurations, initialization strategies, and server-side aggregation bias.
+Personalized federated fine-tuning methods [15, 45, 30] have proposed combining global and local LoRA modules, but often suffer from cross-client interference during aggregation or depend on server-side data.
+Our method introduces a robust and scalable approach to personalized fine-tuning across heterogeneous clients.
+A more comprehensive review of related work is provided in appendix C.
+6 Conclusion
+In this work, we introduced FedALT, a personalized federated LoRA fine-tuning framework that departs from conventional FedAvg paradigms.
+Our key innovations include a decoupled training scheme with Individual and frozen Rest-of-World (RoW) LoRA components that mitigates harmful cross-client interference in heterogeneous settings.
+Additionally, we introduce an adaptive mixer inspired by Mixture-of-Experts that dynamically balances local and global information on an input-specific basis.
+Through comprehensive evaluation across diverse NLP tasks, FedALT demonstrates superior performance compared to state-of-the-art baselines while maintaining computational efficiency.
+Future research directions include investigating clustering-based approaches for maintaining multiple RoW LoRA components, and extending our framework to multi-modal foundation models.
+Appendix A Datasets
+In this paper, we follow the setup proposed in [45] to construct federated datasets for heterogeneous client settings using the Flan collection [6].
+The Flan benchmarks comprises over 12 NLP task categories, each containing multiple datasets.
+These tasks, derived from diverse contextual domains, inherently involve complex distribution shifts.
+We consider three federated dataset configurations.
+In the High Heterogeneous setting, 8 clients are assigned entirely different NLP tasks, resulting in a total of 8 distinct task types.
+In the Low Heterogeneous setting, 8 clients are distributed across only two tasks—Sentiment Analysis and Natural Language Inference—with all clients working on the same task drawing data from a common source, thereby minimizing heterogeneity.
+The Mild Heterogeneous setting assigns 8 clients to datasets spanning 6 different NLP tasks.
+To better simulate realistic FL environments, where each client has limited access to large-scale data, we apply a downsampling strategy: each client is assigned 600 training samples and 300 testing samples.
+For our extended evaluations with 16 and 24 clients, we maintain the same total dataset size while distributing it across more clients.
+Specifically, in the 16-client configuration, each client receives 300 training samples and 150 testing samples, while in the 24-client setting, each client is allocated 200 training samples and 100 testing samples.
+The specific tasks and datasets used in each configuration are provided in tables 5, 6 and 7.
+While some datasets appear in multiple configurations, the sample selections are made independently and randomly.
+The original Flan tasks exhibit highly diverse formats, making standardized preprocessing and evaluation across LLMs challenging.
+Even within a single dataset (e.g., Sentiment140), sample formats can vary significantly, as illustrated in tables 8 and 9.
+To ensure consistency, we adopt a unified prompt template based on [45], shown in table 10.
+For evaluation, we follow FedDPA [45] and use ROUGE-1 as the metric across all tasks, providing a unified measure of model performance.
+This choice is particularly appropriate as Flan inherently structures NLP tasks in a generative format, where ROUGE-1 effectively captures lexical overlap between generated outputs and reference answers.
+Using ROUGE-1 enables direct comparability across diverse tasks and clients, which is essential for analyzing federated learning dynamics.
+While task-specific metrics might offer more nuanced evaluation for individual tasks, they would introduce incomparability issues that complicate assessment of the overall federated learning performance, making ROUGE-1 a methodologically sound choice.
+Table 5: Tasks and Datasets for High Heterogeneous
+Task Dataset Description Text Classification ag_news_subset News classification into categories.
+Natural Language Inference snli Determine relationships between sentence pairs.
+Reading Comprehension openbookqa Open-book question answering.
+Paraphrase Detection glue_mrpc Identify semantic equivalence in sentences.
+Commonsense Reasoning story_cloze Choose the correct story ending.
+Struct to Text common_gen Generate text from concepts.
+Sentiment Analysis sentiment140 Sentiment analysis of tweets.
+Coreference Resolution definite_pronoun_resolution Resolve ambiguous pronouns.
+Table 6: Tasks and Datasets for Mild Heterogeneous
+Task Dataset Description Struct to Text 1 common_gen Generate text from concepts.
+Struct to Text 2 dart Generate structured text descriptions.
+Coreference Resolution definite_pronoun_resolution Resolve ambiguous pronouns.
+Natural Language Inference 1 snli Determine relationships between sentence pairs.
+Natural Language Inference 2 qnli Answer questions based on context.
+Paraphrase Detection glue_mrpc Identify semantic equivalence in sentences.
+Text Classification ag_news_subset Classify news into categories.
+Sentiment Analysis sst2 Analyze sentiment of text.
+Table 7: Tasks and Datasets for Low Heterogeneous
+Task Dataset Description Sentiment Analysis 1 sst2 Analyze sentiment of short movie reviews.
+Sentiment Analysis 2 sentiment140 Sentiment analysis of tweets.
+Natural Language Inference 1 qnli Answer questions based on context.
+Natural Language Inference 2 snli Determine relationships between sentence pairs.
+Table 8: Examples of Data in Sentiment Analysis Task.
+Data Examples Instruction: @celinieee The crafts thingy is already done.
+Describe the sentiment embodied by this tweet.
+OPTIONS: - negative - positive.
+Output: negative Dataset: sentiment140 Instruction: Write a positive tweet.
+Output: @MandinaM Very well indeed thank you..
+Yourself?
+Just about to start work..
+9 hours to go..
+Dataset: sentiment140
+Table 9: Examples of Data in Struct to Text Task.
+Data Examples Instruction: What are the most important words in the following sentence: pig in a pen on a pig farm Output: farm, pen, pig Dataset: common_gen Instruction: Generate a sentence about this data: Wildwood eatType coffee shop; Wildwood priceRange moderate; Wildwood customer rating 1 out of 5; Wildwood near Ranch Output: The Wildwood coffee shop is an average-priced location that serves a light menu of food and alcohol.
+Located near the Ranch, customers give this location a 1 out of 5 stars.
+Dataset: dart
+Table 10: Prompt Template.
+Template
+Prompt Input Instruction: {instruction} Response:
+Appendix B Training Implementations
+We implement all methods using PyTorch and conduct all experiments on six NVIDIA A100 GPUs.
+For the federated methods, we set the local epochs to 5 and the global rounds to 20.
+For the Local Only baseline, the total number of training rounds is set to 50.
+The learning rate is explored in the range $\{3\times 10^{-2},1\times 10^{-3},3\times 10^{-3},1\times 10^{-4},3\times 10^{-4},1\times 10^{-5}\}$, and the AdamW optimizer is utilized, following [45].
+LoRA modules are inserted into the $q$ and $v$ components of each self-attention layer.
+For LLaMA2 experiments, we set the finetuning batch size to 8 and the inference batch size to 2.
+For Bloom experiments, both finetuning and inference batch sizes are set to 32.
+The rank of the trainable LoRA for each method is primarily set to 8, with additional experiments using ranks 4 and 6 conducted for sensitivity analysis.
+We configure other LoRA hyperparameters with lora_alpha set to 32 and lora_dropout set to 0.05.
+Due to computational constraints, the pre-trained LLaMA2 model is loaded using 8-bit quantization to reduce memory requirements while maintaining performance.
+Appendix C Related Work and Compared Methods
+C.1 Parameter-Efficient Fine-Tuning
+As the size of LLMs continues to grow, full parameter fine-tuning has become increasingly computationally and storage-intensive [13].
+To address this issue, Parameter-Efficient Fine-Tuning (PEFT) methods [9, 11, 14, 25, 24, 23] have been developed to significantly reduce the number of trainable parameters.
+PEFT techniques introduce a limited set of additional trainable parameters to enhance model performance while keeping the majority of pre-trained parameters frozen.
+Among all existing PEFT methods, the most prominent approach is LoRA [16], which employs low-rank matrices to approximate the pre-trained weight matrix, updating only the low-rank components.
+LoRA has gained significant attention recently, becoming a standard method for adapting LLMs like Llama under limited computational resources [22].
+Several works have been proposed to improve LoRA [36, 22, 41, 32, 26].
+For instance, a recent development, Hydra-LoRA [36], employs an asymmetric structure with a shared matrix for all samples and distinct matrices for each intrinsic component, thereby enhancing domain adaptation.
+While these advancements primarily focus on centralized learning settings, our work extends the application of PEFT methods to federated LLM fine-tuning, addressing the unique challenges posed by distributed and heterogeneous client data.
+C.2 Federated Learning
+The foundational work in federated learning, FedAvg [27], addresses privacy and communication efficiency by aggregating local model parameters to train a shared global model.
+Since then, numerous studies [39, 42, 46, 2] have focused on tackling various challenges within FL settings.
+One of the key challenges is data heterogeneity, which makes it difficult to train a single shared global model that performs well across all clients.
+To address this, Personalized Federated Learning(PFL) [35, 8, 7, 10] has emerged as an approach to adapt the global model to the specific needs of each client in the federation.
+Most PFL research has focused on addressing statistical differences in data distributions (e.g., label distribution skew) and has been applied primarily to smaller, standard models.
+In contrast, our work addresses the unique challenges posed by heterogeneous clients with diverse NLP tasks by leveraging LLM fine-tuning within the FL framework.
+C.3 Federated Fine-Tuning
+Several studies [1, 5, 43, 20, 44, 3] have explored federated fine-tuning approaches.
+[20] proposed federated fine-tuning with all parameters updated, while [33] introduced federated fine-tuning with PEFT using prefix-tuning.
+[49] was the first study to apply LoRA in a federated context.
+[1, 5] focused on federated fine-tuning where clients have different LoRA ranks, while [43] proposed strategies for LoRA initialization to achieve better performance.
+Another studies [3, 34] addressed server aggregation bias in LoRA-based federated fine-tuning.
+The most relevant works to ours focus on personalized federated fine-tuning with LoRA [15, 45, 30].
+FedDPA [45] introduced global and local LoRA components but sufferd from interference during global LoRA training and struggled to balance the two phases effectively.
+PF2LoRA [15] employed a bilevel framework to combine a shared adapter with client-specific adapters but faced interference issues during the aggregation of the shared adapter.
+FDLoRA [30] utilized personalized LoRA modules to initialize federated training, combining them with a global LoRA module via adaptive fusion; however, it relied on server-side datasets and was prone to interference during global LoRA aggregation, limiting its effectiveness in heterogeneous settings.
+A detailed overview of the baseline methods used for comparison in our experiments is provided below.
+FedIT [49].
+FedIT integrates LoRA with the FedAvg framework, making it one of the earliest approaches to combine these techniques.
+However, under heterogeneous client data settings, FedIT suffers from harmful cross-client interference during the aggregation step, as highlighted in the motivational experiments.
+This interference often reduces its efficiency and, in some cases, results in worse performance than the Local Only method.
+FFA-LoRA [34].
+FFA-LoRA aims to reduce communication costs by fixing the randomly initialized non-zero $\mathbf{A}$ matrices and fine-tuning only the zero-initialized $\mathbf{B}$ matrices.
+While this approach lowers communication overhead, it impairs the learning capacity of LoRA, as fixed $\mathbf{A}$ matrices limit the model’s ability to adapt to diverse client data, resulting in suboptimal performance.
+FedSA [12].
+FedSA introduces an asymmetric approach to LoRA matrix updates in federated learning, where only the $\mathbf{A}$ matrices are shared with the server for aggregation, while the $\mathbf{B}$ matrices remain local to capture client-specific knowledge.
+However, even the client-specific $\mathbf{B}$ matrices may contain valuable knowledge that could benefit other clients.
+The failure to leverage this knowledge effectively results in reduced overall performance.
+PF2LoRA [15].
+PF2LoRA adopts a two-level low-rank adaptation framework, where a common adapter shared across all clients is combined with client-specific lightweight adapters for personalization.
+This bilevel optimization framework aims to balance shared knowledge and personalization.
+However, the aggregation of the shared adapter during the server aggregation step can reintroduce harmful interference, limiting the method’s effectiveness in highly heterogeneous settings.
+FDLoRA [30].
+FDLoRA follows a two-step process where personalized LoRA modules are first trained locally and then used to initialize federated training to obtain a global LoRA module.
+An adaptive fusion mechanism combines the two LoRA modules during fine-tuning.
+While this method improves personalization, it requires additional server-side datasets for effective global LoRA training.
+Moreover, the aggregation step during global LoRA training remains prone to interference, reducing its robustness in handling client heterogeneity.
+FedDPA [45].
+FedDPA is the most closely related to our proposed method, as it also targets settings where clients possess heterogeneous tasks (not just label non-IID data).
+FedDPA employs both global and local LoRA modules, and enhances personalization by introducing a local fine-tuning step excluded from aggregation.
+However, our proposed FedALT departs from FedDPA in several critical ways: First, in both variants of FedDPA, the global LoRA module is trained independently of the local LoRA modules.
+During global training, the local modules are excluded, leading to persistent interference—similar to FedIT—and the local LoRA modules are never shared across clients.
+In contrast, FedALT introduces the Rest-of-World (RoW) LoRA, which captures knowledge from other clients via aggregation, allowing more effective information sharing.
+Second, FedALT employs a dynamic mixer that enables each client to adaptively balance the contributions of its Individual and RoW LoRA modules on a per-input basis, enabling finer-grained personalization.
+Third, FedDPA separates local and global LoRA training into distinct phases, introducing additional hyperparameters to tune the duration of each phase.
+Short local training reduces personalization; overly long training approximates local-only fine-tuning, diminishing the benefits of FL.
+In contrast, FedALT removes the need for phase separation and additional hyperparameters, simplifying training and improving robustness.
+Appendix D The Algorithm of FedALT
+We summarize the pseudocode of FedALT in Algorithms 1.
+Input: Pretrained model $W_{0}$, number of clients $K$, communication rounds $T$
+1 for each client $k$ do
+2       Initialize Individual LoRA $(A_{k}^{L},B_{k}^{L})$ and Mixer $G_{k}$ randomly;
+3       Initialize RoW LoRA $(A_{k}^{R},B_{k}^{R})\leftarrow 0$;
+4
+5for each round $t=1$ to $T$ do
+       // Client-side training
+6       for each client $k$ in parallel do
+7             Receive RoW LoRA $(A_{k}^{R},B_{k}^{R})$ from server;
+8             for each local epoch do
+9                   Update $(A_{k}^{L},B_{k}^{L})$ and Mixer $G_{k}$ using gradient descent;
+                   // Model output:
+10                   $y=W_{0}x+\alpha_{k}(x)B_{k}^{L}A_{k}^{L}x+(1-\alpha_{k}(x))B_{k}^{R}A_{k}^{R}x$;
+11                   where $\alpha_{k}(x)=\text{softmax}(G_{k}x)$
+12            Upload $(A_{k}^{L},B_{k}^{L})$ to server;
+13
+14
+      // Server-side aggregation
+15       for each client $k$ do
+16             Compute RoW LoRA: $A_{k}^{R}\leftarrow\frac{1}{K-1}\sum_{m\neq k}A_{m}^{L}$, $B_{k}^{R}\leftarrow\frac{1}{K-1}\sum_{m\neq k}B_{m}^{L}$;
+17             Send $(A_{k}^{R},B_{k}^{R})$ to client $k$;
+18
+19
+Algorithm 1 FedALT
+Appendix E Additional Experiment Results
+E.1 Training Workload
+In table 11, we compare the training workload of our proposed FedALT with existing baselines using the Bloom-560M model.
+We report both the percentage of trainable parameters, the average local epoch training time and the inference time.
+FedIT serves as a representative baseline, as most federated methods share similar architecture in terms of parameter count and inference cost.
+Among all methods, FFA-LoRA has the fewest trainable parameters, as it updates only the LoRA $\mathbf{B}$ matrices; however, this results in significantly lower performance.
+In contrast, FedALT delivers substantially better results while maintaining comparable training/inference time and parameter overhead, demonstrating that it achieves efficiency without sacrificing effectiveness.
+To further ensure that FedALT’s performance gains are not simply due to having more parameters, we conduct two controlled comparisons, as shown in table 12.
+In RoW-Update (rank=8), both Individual and RoW LoRA modules (each with rank 8) are trained jointly with a dynamic mixer.
+In Fixed Weight (rank=16), only the Individual LoRA (rank 16) is trained, while the frozen RoW LoRA is combined with a fixed mixing weight $\alpha=0.5$.
+Although both configurations involve more trainable parameters than FedALT, they yield lower performance.
+These comparisons validate two core design choices in FedALT: the effectiveness of decoupled training and the benefit of dynamic mixing for personalized adaptation.
+Table 11: Comparison of training workload. FedALT uses a comparable number of LoRA parameters to other baselines.
+Methods Trainable Parameters Avg Local Epoch Time Inference Time FedIT 0.0622% 6.66 s 8.89 s FFA-LoRA 0.0311% 6.56 s 8.77 s FedALT 0.0661% 7.05 s 8.94 s
+Table 12: Comparison of trainable parameters and performance. FedALT achieves the highest performance while using fewer parameters.
+Methods Trainable Parameters Avg Performance
+RoW-Update (rank=8) 0.1283% 65.75
+Fixed Weight (rank=16) 0.1244% 64.59
+FedALT 0.0661% 67.55
+E.2 Per-Client Performance
+Due to space limitations, we reported only the average results for the Low Heterogeneous and Mild Heterogeneous settings in the main paper.
+Here, we present the detailed per-client performance in tables 13, 14, 15 and 16.
+Table 13: Performance on mild heterogeneous (LLaMA2-7B). Evaluation across tasks in Table 6.
+Methods LLaMA2-7B Average
+Struct 1 Struct 2 Coref. NLI 1 NLI 2 Para. Text Cls. Sentiment
 Local Only 52.98 57.35 74.08 66.10 69.02 78.67 64.97 70.16 66.67
@@
 FedALT 51.73 60.72 81.91 74.51 71.12 80.33 70.28 73.22 70.48
-Table 2: Performance comparison with baseline methods in the second federated dataset setting. The average column represents the overall performance across these tasks. See details in Section 4.1.
-4.1 Experiments Results
-Overall Performance Comparisons.
-We compare the performance of the proposed method, FedALT, with other baseline methods based on each client’s performance on its specific test dataset, aligned with its training dataset.
-As shown in Table 1, which reports the results for the first federated setting, our proposed method achieves the best overall performance compared to all baselines, with the average score exceeding that of the best-performing baseline, FDLoRA, by 2.13.
-FedALT consistently delivers superior performance across most tasks, outperforming the baselines by at least 3.27 on the Natural Language Inference task and 3.50 on the Reading Comprehension task.
-We find that on the Struct-to-Text, the Local Only method achieves the highest performance.
-This is likely due to the unique nature of this task, making it difficult to benefit from shared global knowledge.
-Furthermore, we observe that FedIT underperforms compared to Local Only on four client tasks, indicating the presence of harmful interference during aggregation.
-Personalized federated LoRA fine-tuning methods (i.e., FedDPA, PF2LoRA, FDLoRA) also fall short compared to FedALT, likely because they fail to effectively isolate harmful interference during the server aggregation step.
-Despite their designs for local/global LoRA interaction, these methods update LoRA modules using aggregated global information, which may carry conflicting information from other clients.
-To further validate the effectiveness of FedALT, Table 2 presents the results for the second federated setting, involving a different mix of NLP tasks, clients, and task types.
-The results align with those in the first setting, with FedALT outperforming all baseline methods.
-Notably, FedALT achieves an average score of 70.48, surpassing the baseline method by at least 2.88.
-This consistent performance across diverse settings highlights the robustness and adaptability of FedALT, demonstrating its ability to handle heterogeneous data distributions effectively.
-Additionally, in Table 3, we report the number of trainable parameters and the parameters required for inference for both the baseline methods and our proposed method.
-Due to space limitations, we present FedIT as an example, as the other federated learning-based methods have identical trainable and inference parameters.
-All methods exhibit a similar number of trainable parameters, except for FFA-LoRA, which has the fewest trainable parameters since it only updates the LoRA $\mathbf{B}$ module.
-However, despite its reduced parameter count, FFA-LoRA is significantly outperformed by our proposed method in overall performance.
-Regarding the parameters required for inference, our method involves an additional computation for the RoTW LoRA.
-However, this parameter overhead is negligible compared to the size of the large pre-trained LLM.
-Methods Trainable Parameters Additional Inference Workload
-FedIT 0.0622% 0.0622%
-FFA-LoRA 0.0311% 0.0311%
-FedDPA 0.0622% 0.1245%
-FedALT 0.0661% 0.1283%
-Table 3: Comparison of trainable and inference parameters.
-Note that, except for FFA-LoRA and FedDPA, all other baselines share the same parameter count as FedIT.
-Our method employs a comparable number of LoRA parameters to other baselines, which is negligible compared to the pre-trained LLM.
-See details in Section 4.1.
-Impact of Decoupling LoRA Training.
-In our proposed method, only the Individual LoRA modules are updated locally during client fine-tuning, while the RoTW LoRA remains frozen and is updated solely through server-side aggregation.
-This design aims to mitigate harmful cross-client interference by preventing local updates to the shared global component.
-To validate the effectiveness of this decoupled training approach, we compare FedALT against two alternative methods.
-The first comparison method, referred to as RoTW-Update (rank = 4), fine-tunes both the Individual LoRA and RoTW LoRA locally, with each LoRA having a rank of 4 to maintain an equivalent total number of trainable parameters.
-The second comparison method, RoTW-Update (rank = 8), also fine-tunes both LoRA modules locally but assigns a rank of 8 to each, resulting in a higher total parameter count.
-The results, shown in Table 4, reveal that while the second method outperforms the first due to its larger parameter count, both alternative methods underperform compared to our proposed approach.
-Notably, even with more trainable parameters, the second method fails to match the performance of FedALT.
-Methods Same Trainable Param. Same Inference Param. Avg. Performance
-RoTW-Update (rank = 4) $\checkmark$ $\times$ 63.59
-RoTW-Update (rank = 8) $\times$ $\checkmark$ 65.75
-FedALT 67.55
-Table 4: Impact of decoupling LoRA training. Updating the RoTW LoRA component locally leads to harmful cross-client interference. See details in Section 4.1.
-Figure 3: Impact of Dynamic Weighting via MOE.
-The proposed dynamic personalized mixer mechanism surpasses both fixed-weight approaches and shared averaged mixers across clients.
-See Section 4.1 for details.
-Impact of Dynamic Weighting via MOE.
-To enable input-specific flexibility, our proposed method adopts a MOE mechanism to dynamically adjust the contributions of the Individual LoRA and RoTW LoRA.
-We conduct experiments to evaluate the efficiency of this dynamic weighting by comparing it to two fixed weighting strategies.
-In the first fixed strategy, we set the contribution weight to a fixed value of $\alpha=0.5$, meaning that the Individual LoRA and RoTW LoRA contribute equally.
-In the second fixed strategy, we set the contribution weight to $\alpha=1/K$, where $K$ is the number of clients.
-This configuration effectively reduces the method to using the averaged LoRA, relying solely on each client’s last-round LoRA modules for the next round’s initialization.
-The results in Figure 3 validate that our proposed dynamic weighting mechanism outperforms both fixed strategies, demonstrating its ability to enhance performance.
-Additionally, our dynamic weighting is implemented through the personal $\mathbf{G}_{k}$ mixer, which is maintained locally by each client and is not aggregated during the training process.
-To further investigate this design choice, we conducted experiments where $\mathbf{G}_{k}$ is averaged across clients.
-The results confirm that retaining a personalized $\mathbf{G}_{k}$ for each client yields optimal performance, emphasizing the importance of input-specific and client-specific adjustments.
-5 Related Work
-5.1 Parameter-Efficient Fine-Tuning
-As the size of LLMs continues to grow, full parameter fine-tuning has become increasingly computationally and storage-intensive (Hadi et al., 2023).
-To address this issue, Parameter-Efficient Fine-Tuning (PEFT) methods (Ding et al., 2023; Fu et al., 2023; Han et al., 2024; Liu et al., 2022; Li & Liang, 2021; Lester et al., 2021) have been developed to significantly reduce the number of trainable parameters.
-PEFT techniques introduce a limited set of additional trainable parameters to enhance model performance while keeping the majority of pre-trained parameters frozen.
-Among all existing PEFT methods, the most prominent approach is LoRA (Hu et al., 2021), which employs low-rank matrices to approximate the pre-trained weight matrix, updating only the low-rank components.
-LoRA has gained significant attention recently, becoming a standard method for adapting LLMs like Llama under limited computational resources (Lermen et al., 2023).
-Several works have been proposed to improve LoRA (Tian et al., 2024; Lermen et al., 2023; Wang et al., 2023; Sheng et al., 2023; Liu et al., 2023).
-For instance, a recent development, Hydra-LoRA (Tian et al., 2024), employs an asymmetric structure with a shared matrix for all samples and distinct matrices for each intrinsic component, thereby enhancing domain adaptation.
-While these advancements primarily focus on centralized learning settings, our work extends the application of PEFT methods to federated LLM fine-tuning, addressing the unique challenges posed by distributed and heterogeneous client data.
-5.2 Federated Learning
-The foundational work in federated learning, FedAvg (McMahan et al., 2017), addresses privacy and communication efficiency by aggregating local model parameters to train a shared global model.
-Since then, numerous studies (Wang et al., 2024a, c; Yang et al., 2024b; Bian et al., 2024a) have focused on tackling various challenges within FL settings.
-One of the key challenges is data heterogeneity, which makes it difficult to train a single shared global model that performs well across all clients.
-To address this, Personalized Federated Learning(PFL) (Tan et al., 2022; Deng et al., 2020; Collins et al., 2021; Fallah et al., 2020) has emerged as an approach to adapt the global model to the specific needs of each client in the federation.
-Most PFL research has focused on addressing statistical differences in data distributions (e.g., label distribution skew) and has been applied primarily to smaller, standard models.
-In contrast, our work addresses the unique challenges posed by heterogeneous clients with diverse NLP tasks by leveraging LLM fine-tuning within the FL framework.
-5.3 Federated Fine-Tuning
-Several studies (Bai et al., 2024; Cho et al., 2024; Wang et al., 2024d; Kuang et al., 2024; Wu et al., 2024; Bian et al., 2024b) have explored federated fine-tuning approaches.
-(Kuang et al., 2024) proposed federated fine-tuning with all parameters updated, while (Sun et al., 2022) introduced federated fine-tuning with PEFT using prefix-tuning.
-(Zhang et al., 2024) was the first study to apply LoRA in a federated context.
-(Bai et al., 2024; Cho et al., 2024) focused on federated fine-tuning where clients have different LoRA ranks, while (Wang et al., 2024d) proposed strategies for LoRA initialization to achieve better performance.
-Another studies (Bian et al., 2024b; Sun et al., 2024) addressed server aggregation bias in LoRA-based federated fine-tuning.
-The most relevant works to ours focus on personalized federated fine-tuning with LoRA (Anonymous, 2025; Yang et al., 2024a; Qi et al., 2024).
-FedDPA (Yang et al., 2024a) introduced global and local LoRA components but sufferd from interference during global LoRA training and struggled to balance the two phases effectively.
-PF2LoRA (Anonymous, 2025) employed a bilevel framework to combine a shared adapter with client-specific adapters but faced interference issues during the aggregation of the shared adapter.
-FDLoRA (Qi et al., 2024) utilized personalized LoRA modules to initialize federated training, combining them with a global LoRA module via adaptive fusion; however, it relied on server-side datasets and was prone to interference during global LoRA aggregation, limiting its effectiveness in heterogeneous settings.
-6 Conclusion
-In this work, We proposed FedALT, a personalized federated LoRA fine-tuning framework that departs from the FedAvg paradigm by prioritizing local adaptation while integrating global knowledge via a Rest-of-the-World (RoTW) LoRA component.
-An adaptive mixer dynamically balances local and global information, mitigating cross-client interference.
-Experiments across NLP tasks demonstrate FedALT’s superiority in heterogeneous settings with minimal overhead.
-Future work will focus on a theoretical analysis of FedALT’s convergence.
-Appendix A Datasets
-In this paper, we primarily follow the setup of (Yang et al., 2024a) to construct federated datasets for heterogeneous client settings using the Flan datasets (Chung et al., 2024).
-The Flan collection consists of over 12 NLP tasks, each containing multiple datasets.
-These tasks, generated from diverse contextual factors, inherently experience various complex distribution shifts.
-We consider two federated dataset configurations.
-For the first configuration, we assume an extreme client heterogeneity setting, where 8 clients are assigned completely different NLP tasks (a total of 8 distinct tasks are considered).
-For the second configuration, we consider a relatively moderate setting, where some clients share the same NLP task but with data originating from different datasets.
-In this setting, 8 clients with 6 distinct NLP tasks are included.
-To simulate a realistic FL environment, where each client has limited access to large datasets, we employ a downsampling strategy, assigning each client 600 training samples and 300 testing samples.
-The specific tasks and datasets for each federated dataset configuration are detailed in Tables 5 and 6.
-Note that while there is some overlap in the datasets used for federated setting 1 and federated setting 2, the samples within each configuration are randomly selected, ensuring independence between the two settings.
-The original NLP tasks in Flan exhibit diverse formats, making it challenging to standardize processing for LLMs and to select a consistent evaluation metric.
-Even within a single dataset (e.g., Sentiment140), samples can vary significantly in format, as illustrated in Tables 7 and 8.
-To facilitate uniform processing and evaluation, we adopt a simple prompt template following (Yang et al., 2024a), as shown in Table 9.
-ROUGE-1 is chosen as the evaluation metric to assess performance across all tasks.
-Task Dataset Description
-Text Classification ag_news_subset News classification into categories.
-Natural Language Inference snli Determine relationships between sentence pairs.
-Reading Comprehension openbookqa Open-book question answering.
-Paraphrase Detection glue_mrpc Identify semantic equivalence in sentences.
-Commonsense Reasoning story_cloze Choose the correct story ending.
-Struct to Text common_gen Generate text from concepts.
-Sentiment Analysis sentiment140 Sentiment analysis of tweets.
-Coreference Resolution definite_pronoun_resolution Resolve ambiguous pronouns.
-Table 5: Tasks and Datasets for Federated Dataset 1
-Task Dataset Description
-Struct to Text 1 common_gen Generate text from concepts.
-Struct to Text 2 dart Generate structured text descriptions.
-Coreference Resolution definite_pronoun_resolution Resolve ambiguous pronouns.
-Natural Language Inference 1 snli Determine relationships between sentence pairs.
-Natural Language Inference 2 qnli Answer questions based on context.
-Paraphrase Detection glue_mrpc Identify semantic equivalence in sentences.
-Text Classification ag_news_subset Classify news into categories.
-Sentiment Analysis sst2 Analyze sentiment of text.
-Table 6: Tasks and Datasets for Federated Dataset 2
-Data Examples
-Instruction: @celinieee The crafts thingy is already done. Describe the sentiment embodied by this tweet. OPTIONS: - negative - positive.
-Output: negative
-Dataset: sentiment140
-Instruction: Write a positive tweet.
-Output: @MandinaM Very well indeed thank you.. Yourself? Just about to start work.. 9 hours to go..
-Dataset: sentiment140
-Table 7: Examples of Data in Sentiment Analysis Task.
-Data Examples
-Instruction: What are the most important words in the following sentence: pig in a pen on a pig farm
-Output: farm, pen, pig
-Dataset: common_gen
-Instruction: Generate a sentence about this data: Wildwood eatType coffee shop; Wildwood priceRange moderate; Wildwood customer rating 1 out of 5; Wildwood near Ranch
-Output: The Wildwood coffee shop is an average-priced location that serves a light menu of food and alcohol. Located near the Ranch, customers give this location a 1 out of 5 stars.
-Dataset: dart
-Table 8: Examples of Data in Struct to Text Task.
-Template
-Prompt Input Instruction: {instruction} Response:
-Table 9: Prompt Template.
-Appendix B Training Implementations
-We implement all methods using PyTorch and conduct all experiments on six NVIDIA A100 GPUs.
-For the federated methods, we set the local epochs to 5 and the global rounds to 20.
-For the Local Only baseline, the total number of training rounds is set to 50.
-The learning rate is set to $3\times 10^{-4}$, and the AdamW optimizer is utilized, following (Yang et al., 2024a).
-LoRA modules are inserted into the $q$ and $v$ components of each self-attention layer in LLaMA2.
-The batch size is set to 8, and the rank of the trainable LoRA for each method is fixed at 8.
-Additionally, we set lora_alpha to 32 and lora_dropout to 0.05.
-Due to computational limitations, the pre-trained LLaMA2 model is loaded in 8-bit precision.
-Appendix C Methods
-Local Only.
-The Local Only method represents a setting where each client independently trains its own model without sharing any model updates with a server.
-This method serves as a baseline to assess the impact of federated learning methods on model performance.
-FedIT (Zhang et al., 2024).
-FedIT integrates LoRA with the FedAvg framework, making it one of the earliest approaches to combine these techniques.
-However, under heterogeneous client data settings, FedIT suffers from harmful cross-client interference during the aggregation step, as highlighted in the motivational experiments.
-This interference often reduces its efficiency and, in some cases, results in worse performance than the Local Only method.
-FFA-LoRA (Sun et al., 2024).
-FFA-LoRA aims to reduce communication costs by fixing the randomly initialized non-zero $\mathbf{A}$ matrices and fine-tuning only the zero-initialized $\mathbf{B}$ matrices.
-While this approach lowers communication overhead, it impairs the learning capacity of LoRA, as fixed $\mathbf{A}$ matrices limit the model’s ability to adapt to diverse client data, resulting in suboptimal performance.
-FedSA (Guo et al., 2024).
-FedSA introduces an asymmetric approach to LoRA matrix updates in federated learning, where only the $\mathbf{A}$ matrices are shared with the server for aggregation, while the $\mathbf{B}$ matrices remain local to capture client-specific knowledge.
-However, even the client-specific $\mathbf{B}$ matrices may contain valuable knowledge that could benefit other clients.
-The failure to leverage this knowledge effectively results in reduced overall performance.
-FedDPA (Yang et al., 2024a).
-FedDPA is a state-of-the-art personalized federated fine-tuning method that introduces two distinct LoRA components: global LoRA and local LoRA.
-The global LoRA is trained using FedAvg, while the local LoRA is independently fine-tuned on client-specific data.
-Despite its personalization capabilities, FedDPA suffers from interference during the global LoRA training stage.
-Additionally, its separation of local and global LoRA fine-tuning into distinct phases complicates the determination of optimal training durations for each phase.
-PF2LoRA (Anonymous, 2025).
-PF2LoRA adopts a two-level low-rank adaptation framework, where a common adapter shared across all clients is combined with client-specific lightweight adapters for personalization.
-This bilevel optimization framework aims to balance shared knowledge and personalization.
-However, similar to FedDPA, the aggregation of the shared adapter during the server aggregation step can reintroduce harmful interference, limiting the method’s effectiveness in highly heterogeneous settings.
-FDLoRA (Qi et al., 2024).
-FDLoRA follows a two-step process where personalized LoRA modules are first trained locally and then used to initialize federated training to obtain a global LoRA module.
-An adaptive fusion mechanism combines the two LoRA modules during fine-tuning.
-While this method improves personalization, it requires additional server-side datasets for effective global LoRA training.
-Moreover, the aggregation step during global LoRA training remains prone to interference, reducing its robustness in handling client heterogeneity.
-Appendix D Impact of RoTW LoRA
-To evaluate the impact of using the RoTW LoRA instead of the global average of all Local LoRAs (which includes each client’s Individual LoRA), we conduct an experiment.
-As shown in Figure 4, leveraging the RoTW LoRA significantly improves performance compared to using the global average.
-This result underscores the importance of isolating client-specific information from the aggregated global knowledge.
-By decoupling the RoTW LoRA from client updates, our method effectively mitigates cross-client interference, enabling better personalization.
-Figure 4: Impact of RoTW LoRA. Results indicate that isolating client-specific information is more effective than directly using globally averaged information. See details in Appendix D.
-Appendix E Discussions
-E.1 Limitations of FedALT in Traditional FL
-In FedALT, a separate RoTW LoRA is maintained alongside each client’s Individual LoRA component.
-The RoTW LoRA is not trained during each round but remains fixed, which slightly increases the forward/inference workload.
+Table 14: Performance on mild heterogeneous (Bloom-560M). Evaluation across tasks in Table 6.
+Methods Bloom-560M Average
+Struct 1 Struct 2 Coref. NLI 1 NLI 2 Para. Text Cls. Sentiment
+Local 40.47 48.10 34.48 42.33 52.58 70.26 47.65 65.82 50.21
+FedIT 35.38 41.20 38.86 39.23 49.16 71.67 51.75 65.00 49.03
+FFA-LoRA 36.16 39.40 36.57 36.22 47.19 68.67 50.45 63.89 47.32
+FedSA 35.98 44.71 38.68 47.68 55.92 70.54 58.02 68.84 52.55
+FedDPA 38.56 42.54 39.23 46.69 49.21 69.00 62.35 68.92 52.06
+PF2LoRA 39.65 42.29 40.36 45.85 53.07 70.40 57.43 66.67 51.97
+FDLoRA 40.66 44.62 38.91 43.54 51.17 70.33 54.13 68.00 51.42
+FedALT 41.10 48.64 39.80 48.14 56.63 71.27 64.84 68.67 54.89
+Table 15: Performance on low heterogeneous (LLaMA2-7B). Evaluation across 8 clients split by task-dataset pairs. SA1–SA2 use sst2, SA3–SA4 use sentiment140, NLI1–NLI2 use qnli, and NLI3–NLI4 use snli.
+Methods SA1 SA2 SA3 SA4 NLI1 NLI2 NLI3 NLI4 Average
+Local 65.95 63.95 63.57 66.34 45.95 47.45 66.93 64.60 60.59
+FedIT 67.25 66.82 65.85 64.79 46.53 49.48 68.53 68.65 62.24
+FFA-LoRA 66.42 61.84 62.96 65.14 46.83 49.23 67.93 68.16 61.07
+FedSA 71.85 64.31 65.45 69.37 46.31 48.24 67.03 70.38 62.87
+FedDPA 67.39 64.74 65.58 63.98 46.89 48.63 69.77 69.47 62.06
+PF2LoRA 71.73 67.83 65.45 70.18 47.19 47.68 69.41 69.03 63.56
+FDLoRA 68.54 65.46 65.93 63.36 47.09 49.19 70.69 69.07 62.42
+FedALT 73.05 67.86 65.91 72.82 47.29 50.01 70.42 69.88 64.66
+Table 16: Performance on low heterogeneous (Bloom-560M). Each dataset is split across two clients. See Table 7 for dataset mapping.
+Methods SA1 SA2 SA3 SA4 NLI1 NLI2 NLI3 NLI4 Average
+Local 44.33 40.45 46.03 51.04 40.64 42.58 62.13 62.03 48.65
+FedIT 46.89 45.16 50.20 54.69 42.49 46.26 64.19 67.29 52.15
+FFA-LoRA 52.18 36.04 52.47 51.91 42.78 45.29 64.50 65.45 51.33
+FedSA 53.55 47.16 51.63 50.47 44.96 45.92 66.36 64.18 53.03
+FedDPA 52.61 47.27 52.03 53.14 43.26 45.56 65.68 65.38 53.12
+PF2LoRA 46.88 43.00 55.24 55.30 39.42 43.39 64.14 61.39 51.10
+FDLoRA 50.66 46.72 49.38 54.07 42.35 45.82 64.63 66.45 52.51
+FedALT 54.06 49.18 53.34 55.78 42.94 45.51 67.57 67.49 54.48
+Appendix F Convergence Analysis
+Since our mixer depends on each sample input, existing convergence analysis frameworks cannot be directly applied.
+In this section, we provide an initial convergence analysis of FedALT under appropriate assumptions.
+We begin by defining the key parameters and establishing necessary conditions.
+While this analysis offers theoretical validation for our approach, we acknowledge that it represents a preliminary investigation into the convergence properties of FedALT.
+More comprehensive theoretical frameworks with potentially stronger guarantees and less restrictive assumptions remain an important direction for future research.
+Nevertheless, our initial analysis provides valuable insights into the convergence behavior and offers a theoretical foundation that complements our empirical findings.
+F.1 Notation and Setup
+Let the trainable parameters for client $k$ at round $t$ be denoted by
+$Z_{k}^{t}=\{\mathbf{A}_{k}^{L},\mathbf{B}_{k}^{L},\mathbf{G}_{k}\}$ (5)
+where $\mathbf{A}_{k}^{L}$ and $\mathbf{B}_{k}^{L}$ represent the Individual LoRA modules and $\mathbf{G}_{k}$ denotes the adaptive mixer.
+The Rest-of-World (RoW) LoRA for client $k$ at round $t$ is computed as:
+$\mathbf{A}_{k}^{R}=\frac{1}{K-1}\sum_{m\neq k}\mathbf{A}_{m}^{L},\quad\mathbf{B}_{k}^{R}=\frac{1}{K-1}\sum_{m\neq k}\mathbf{B}_{m}^{L}$ (6)
+We denote the loss function for client $k$ as $L_{k}(Z_{k}|\mathbf{A}_{k}^{R},\mathbf{B}_{k}^{R})$, which measures the performance on client $k$’s local dataset given its trainable parameters and the fixed RoW LoRA components.
+F.2 Assumptions
+Assumption F.1 (Smoothness).
+For each client $k$, the loss function $L_{k}$ is $\mu$-smooth with respect to $Z_{k}$, i.e., for any $Z_{k}^{1},Z_{k}^{2}$:
+$\|\nabla L_{k}(Z_{k}^{1})-\nabla L_{k}(Z_{k}^{2})\|\leq\mu\|Z_{k}^{1}-Z_{k}^{2}\|$ (7)
+Assumption F.2 (Strong Convexity).
+For each client $k$, the loss function $L_{k}$ is $\lambda$-strongly convex with respect to $Z_{k}$, i.e., for any $Z_{k}^{1},Z_{k}^{2}$:
+$L_{k}(Z_{k}^{2})\geq L_{k}(Z_{k}^{1})+\langle\nabla L_{k}(Z_{k}^{1}),Z_{k}^{2}-Z_{k}^{1}\rangle+\frac{\lambda}{2}\|Z_{k}^{2}-Z_{k}^{1}\|^{2}$ (8)
+Assumption F.3 (Bounded RoW Sensitivity).
+The optimal parameters for client $k$ are sensitive to changes in the RoW parameters with a bounded Lipschitz constant.
+Specifically, for any two sets of RoW parameters $(\mathbf{A}_{k}^{R,1},\mathbf{B}_{k}^{R,1})$ and $(\mathbf{A}_{k}^{R,2},\mathbf{B}_{k}^{R,2})$:
+$\|Z_{k}^{*}(\mathbf{A}_{k}^{R,1},\mathbf{B}_{k}^{R,1})-Z_{k}^{*}(\mathbf{A}_{k}^{R,2},\mathbf{B}_{k}^{R,2})\|\leq\beta\|(\mathbf{A}_{k}^{R,1},\mathbf{B}_{k}^{R,1})-(\mathbf{A}_{k}^{R,2},\mathbf{B}_{k}^{R,2})\|$ (9)
+where $Z_{k}^{*}(\mathbf{A}_{k}^{R},\mathbf{B}_{k}^{R})=\arg\min_{Z_{k}}L_{k}(Z_{k}|\mathbf{A}_{k}^{R},\mathbf{B}_{k}^{R})$ represents the optimal parameters for client $k$ given fixed RoW parameters.
+The first two assumptions are standard and widely used in FL proof.
+The third assumption holds because the RoW LoRA component remains frozen during local training, providing a stable reference point that limits how much client parameters can drift between rounds based on changes in other clients’ parameters.
+F.3 Convergence Theorem
+Theorem F.4 (Convergence of FedALT).
+With the assumptions, we can derive:
+$\|Z_{k}^{t+1}-Z_{k}^{t}\|\leq\frac{2\epsilon}{1-\beta}+\beta^{t}\max_{k}\|Z_{k}^{1}-Z_{k}^{0}\|$
+If $\beta<1$ and the local training at each round converges to a neighborhood of the optimal solution, then the sequence of trainable parameters $\{Z_{k}^{t}\}$ generated by FedALT converges to a stable point for each client $k$.
+Proof.
+We analyze the convergence in two steps:
+  1. 1.
+    Local convergence within each round
+  2. 2.
+    Global convergence across rounds
+Step 1: Local convergence within each round
+At round $t$, client $k$ optimizes its parameters $Z_{k}$ with fixed RoW parameters $(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$.
+Let $Z_{k}^{t,s}$ represent the parameters after $s$ steps of local training.
+We now analyze the convergence of this local optimization process using Assumptions 1 and 2.
+The gradient descent update rule at step $s$ is given by:
+$Z_{k}^{t,s+1}=Z_{k}^{t,s}-\eta\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ (10)
+From Assumption 1 (smoothness), we can have:
+$L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ $\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})+\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{t,s+1}-Z_{k}^{t,s}\rangle$
+$\quad+\frac{\mu}{2}\|Z_{k}^{t,s+1}-Z_{k}^{t,s}\|^{2}$ (11)
+Substituting the update rule:
+$L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ $\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
+$\quad+\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),-\eta\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\rangle$
+$\quad+\frac{\mu}{2}\|-\eta\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (12)
+$=L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-\eta\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$
+$\quad+\frac{\mu\eta^{2}}{2}\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (13)
+$=L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-\eta(1-\frac{\mu\eta}{2})\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (14)
+Let $Z_{k}^{*}=Z_{k}^{*}(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ denote the optimal parameters that minimize the loss function for client $k$ given the fixed RoW parameters.
+From Assumption 2 (strong convexity), we have:
+$L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
+$\geq$ $L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})+\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{*}-Z_{k}^{t,s}\rangle+\frac{\lambda}{2}\|Z_{k}^{*}-Z_{k}^{t,s}\|^{2}$ (15)
+Rearranging:
+$\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{t,s}-Z_{k}^{*}\rangle$
+$\geq$ $L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})+\frac{\lambda}{2}\|Z_{k}^{*}-Z_{k}^{t,s}\|^{2}$ (16)
+By Cauchy-Schwarz inequality:
+$\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|\cdot\|Z_{k}^{t,s}-Z_{k}^{*}\|\geq\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{t,s}-Z_{k}^{*}\rangle$ (17)
+Combining with the previous inequality:
+$\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|\cdot\|Z_{k}^{t,s}-Z_{k}^{*}\|$
+$\geq$ $L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})+\frac{\lambda}{2}\|Z_{k}^{*}-Z_{k}^{t,s}\|^{2}$ (18)
+From the properties of strongly convex functions, we can derive:
+$\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}\geq 2\lambda(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}))$ (19)
+Substituting this into our earlier expression for the loss after one update:
+$L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
+$\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-\eta(1-\frac{\mu\eta}{2})\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (20)
+$\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
+$-2\eta(1-\frac{\mu\eta}{2})\lambda(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}))$ (21)
+$=(1-2\eta\lambda(1-\frac{\mu\eta}{2}))(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})),$ (22)
+where the learning rate $\eta<\frac{2}{\mu}$.
+From Assumption 2 (strong convexity), we have:
+$\frac{\lambda}{2}\|Z_{k}^{t,s}-Z_{k}^{*}\|^{2}\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ (23)
+And from Assumption 1 (smoothness):
+$L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\leq\frac{\mu}{2}\|Z_{k}^{t,s}-Z_{k}^{*}\|^{2}$ (24)
+Combining these inequalities with our convergence rate:
+$\frac{\lambda}{2}\|Z_{k}^{t,s+1}-Z_{k}^{*}\|^{2}$ $\leq L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ (25)
+$\leq(1-2\eta\lambda(1-\frac{\mu\eta}{2}))(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}))$ (26)
+$\leq(1-2\eta\lambda(1-\frac{\mu\eta}{2}))\frac{\mu}{2}\|Z_{k}^{t,s}-Z_{k}^{*}\|^{2}$ (27)
+Simplifying:
+$\|Z_{k}^{t,s+1}-Z_{k}^{*}\|^{2}\leq(1-2\eta\lambda(1-\frac{\mu\eta}{2}))\frac{\mu}{\lambda}\|Z_{k}^{t,s}-Z_{k}^{*}\|^{2}$ (28)
+Denote $\rho=(1-2\eta\lambda(1-\frac{\mu\eta}{2}))\frac{\mu}{\lambda}$, when we set the learning rate be $\frac{1}{\mu}$, we can make $\rho$ in the range of 0 to 1.
+By applying this inequality recursively for $s$ steps:
+$\|Z_{k}^{t,s}-Z_{k}^{*}\|^{2}\leq\rho^{s}\|Z_{k}^{t,0}-Z_{k}^{*}\|^{2}$ (29)
+Taking the square root:
+$\|Z_{k}^{t,s}-Z_{k}^{*}(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|\leq\rho^{s/2}\|Z_{k}^{t,0}-Z_{k}^{*}(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|$ (30)
+Since $\rho<1$, we have $\lim_{s\to\infty}\rho^{s}=0$. Therefore, with a sufficient number of local training steps $S$, we can ensure:
+$\|Z_{k}^{t+1}-Z_{k}^{*}(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|\leq\epsilon$ (31)
+where $Z_{k}^{t+1}=Z_{k}^{t,S}$ and $\epsilon$ can be made arbitrarily small by increasing $S$.
+Step 2: Global convergence across rounds
+Consider the difference between parameters at consecutive rounds:
+$\|Z_{k}^{t+1}-Z_{k}^{t}\|$ $\leq\|Z_{k}^{t+1}-Z_{k}^{*}(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|$ (32)
+$+\|Z_{k}^{*}(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-Z_{k}^{*}(\mathbf{A}_{k}^{R,t-1},\mathbf{B}_{k}^{R,t-1})\|$ (33)
+$+\|Z_{k}^{*}(\mathbf{A}_{k}^{R,t-1},\mathbf{B}_{k}^{R,t-1})-Z_{k}^{t}\|$ (34)
+By Step 1 and Assumption 3:
+$\|Z_{k}^{t+1}-Z_{k}^{t}\|\leq 2\epsilon+\beta\|(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-(\mathbf{A}_{k}^{R,t-1},\mathbf{B}_{k}^{R,t-1})\|$ (35)
+From the definition of RoW LoRA:
+$\|(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-(\mathbf{A}_{k}^{R,t-1},\mathbf{B}_{k}^{R,t-1})\|$ $=\frac{1}{K-1}\|\sum_{m\neq k}(Z_{m}^{t}-Z_{m}^{t-1})\|$ (36)
+$\leq\frac{1}{K-1}\sum_{m\neq k}\|Z_{m}^{t}-Z_{m}^{t-1}\|$ (37)
+Therefore:
+$\|Z_{k}^{t+1}-Z_{k}^{t}\|\leq 2\epsilon+\frac{\beta}{K-1}\sum_{m\neq k}\|Z_{m}^{t}-Z_{m}^{t-1}\|$ (38)
+Let $\Delta^{t}=\max_{k}\|Z_{k}^{t+1}-Z_{k}^{t}\|$. Then:
+$\|Z_{k}^{t+1}-Z_{k}^{t}\|\leq 2\epsilon+\frac{\beta(K-1)}{K-1}\Delta^{t-1}=2\epsilon+\beta\Delta^{t-1}$ (39)
+Taking the maximum over all clients:
+$\Delta^{t}\leq 2\epsilon+\beta\Delta^{t-1}$ (40)
+When $\beta<1$, this is a contraction, and by iterating:
+$\Delta^{t}\leq 2\epsilon\sum_{i=0}^{t-1}\beta^{i}+\beta^{t}\Delta^{0}\leq\frac{2\epsilon}{1-\beta}+\beta^{t}\Delta^{0}$ (41)
+As $t\rightarrow\infty$, $\Delta^{t}\rightarrow\frac{2\epsilon}{1-\beta}$, which can be made arbitrarily small by reducing $\epsilon$ (increasing local training iterations).
+When this condition is satisfied, FedALT converges to a stable solution within a neighborhood determined by the local training precision.
+Appendix G Limitations of FedALT in Traditional FL
+In FedALT, a separate RoW LoRA is maintained alongside each client’s Individual LoRA component.
+The RoW LoRA is not trained during each round but remains fixed, which slightly increases the forward/inference workload.
 However, we emphasize that FedALT is particularly well-suited for federated LoRA fine-tuning.
-Compared to the pre-trained model (e.g., LLaMA), the additional forward computation introduced by the RoTW LoRA is minimal and can be largely neglected.
-However, applying the Individual and RoTW structure to traditional FL, where the full model is updated rather than a lightweight adapter, may not be practical.
-In such a setting, maintaining an additional RoTW model for inference would effectively double the forward computation workload, making it computationally inefficient.
+Compared to the pre-trained model (e.g., LLaMA), the additional forward computation introduced by the RoW LoRA is minimal and can be largely neglected.
+However, applying the Individual and RoW structure to traditional FL, where the full model is updated rather than a lightweight adapter, may not be practical.
+In such a setting, maintaining an additional RoW model for inference would effectively double the forward computation workload, making it computationally inefficient.
 Therefore, our method is specifically designed for federated fine-tuning scenarios, where the overhead remains manageable and the benefits of interference isolation are maximized.
-E.2 Comparison with FedDPA (Yang et al., 2024a)
-FedDPA also introduces two LoRA components, referred to as global LoRA and local LoRA.
-By incorporating an additional local fine-tuning step excluded from the aggregation process, FedDPA aims to enhance personalization for individual clients.
-However, our proposed method, FedALT, differs from FedDPA in several key aspects: First, in both versions of FedDPA, global LoRA training is independent of local LoRA training.
-During global training, the local LoRA components are excluded, meaning that interference persists during the global training process, similar to FedIT.
-Furthermore, because the global LoRA training in FedDPA is independent of local fine-tuning, the local LoRA is never shared with other clients.
-This is in stark contrast to our method, where the RoTW LoRA captures useful training information from all other clients.
-Additionally, the dynamic $\alpha$ weight in FedALT allows each client to individually determine the contribution of its Individual LoRA and Rest-of-World LoRA on the per-input basis, enabling a more personalized and effective model.
-Second, FedDPA separates local and global LoRA fine-tuning into distinct steps, which introduces challenges in determining the duration of each step.
-If the local fine-tuning step is too short, the personalization effect is limited; if it is too long, the process approaches independent local fine-tuning, thereby negating the potential benefits of FL.
-Such additional hyperparameters can significantly affect performance but are often difficult to tune.
-In contrast, our proposed method eliminates the need to separate local and global fine-tuning phases and has no additional hyperparameters, simplifying the process and avoiding these timing-related challenges.

### v3

@@
 Through extensive experiments on NLP benchmarks, we demonstrate that FedALT significantly outperforms state-of-the-art personalized federated LoRA fine-tuning methods, achieving superior local adaptation without sacrificing computational efficiency.
-1 Introduction
+Introduction
 Figure 1: Illustration of FedALT.
@@
 The adaptive mixer dynamically combines the RoW LoRA and Individual LoRA.
-See details in section 3.
-Large language models (LLMs) [19, 4, 31, 37, 50, 48, 38] have demonstrated exceptional capabilities in language understanding and generation, enabling a wide range of applications.
-Adapting pretrained LLMs to specialized domains or further enhancing their performance requires fine-tuning with task-specific data [29].
+Large language models (LLMs) (kenton2019bert; brown2020language; raffel2020exploring; touvron2023llama; zhou2024comprehensive; zeng2022glm; touvron2023llama2) have demonstrated exceptional capabilities in language understanding and generation, enabling a wide range of applications.
+Adapting pretrained LLMs to specialized domains or further enhancing their performance requires fine-tuning with task-specific data (minaee2024large).
 However, full fine-tuning, which involves updating all model parameters, is computationally expensive and often impractical for real-world deployment.
-To address this limitation, several parameter-efficient fine-tuning (PEFT) [9, 11, 14, 25, 24, 23] methods have been proposed, with low-rank adaptation (LoRA) [16] being one of the most widely used.
+To address this limitation, several parameter-efficient fine-tuning (PEFT) (ding2023parameter; fu2023effectiveness; liu2022few; li2021prefix; lester2021power) methods have been proposed, with low-rank adaptation (LoRA) (hu2021lora) being one of the most widely used.
 LoRA reduces the number of trainable parameters by integrating low-rank matrices into the model, significantly reducing computational costs.
 Fine-tuning LLMs necessitates substantial data volumes for adaptation to downstream tasks; however, this requisite data often resides across multiple institutions where privacy regulations and security protocols prohibit direct data sharing.
-Federated Learning (FL) [27] facilitates collaborative fine-tuning without raw data exposure, establishing itself as an essential paradigm for privacy-preserving adaptation.
-Nonetheless, existing federated LoRA fine-tuning methodologies [49, 34, 43, 12, 3] predominantly presuppose a singular global model, neglecting the inherent data heterogeneity among participating clients.
+Federated Learning (FL) facilitates collaborative fine-tuning without raw data exposure, establishing itself as an essential paradigm for privacy-preserving adaptation (bian2025survey).
+Nonetheless, existing federated LoRA fine-tuning methodologies (zhang2024towards; sun2024improving; wang2024flora; Bian_2025_ICCV) predominantly presuppose a singular global model, neglecting the inherent data heterogeneity among participating clients.
 Such a homogeneous modeling approach frequently yields suboptimal performance and introduces fairness disparities, particularly in real-world deployments where client data exhibits significant variance in volume, domain, and distributional characteristics.
-While limited research on personalized federated LoRA fine-tuning [15, 30] exists, these approaches primarily address label heterogeneity cases.
-However, in realistic LLM fine-tuning contexts, clients’ learning objectives naturally diverge due to diverse business imperatives [47].
-Personalized federated LoRA fine-tuning addressing more realistic task-heterogeneous scenarios [45] remains insufficiently investigated in the current literature.
-Federated Averaging (FedAvg) [27] is perhaps the most widely used FL framework, serving as the foundation for numerous follow-up works that introduce various optimizations.
+While limited research on personalized federated LoRA fine-tuning (hao2025personalized; qi2024fdlora) exists, these approaches primarily address label heterogeneity cases.
+However, in realistic LLM fine-tuning contexts, clients’ learning objectives naturally diverge due to diverse business imperatives (yao2022benchmark).
+Personalized federated LoRA fine-tuning addressing more realistic task-heterogeneous scenarios (yang2024dual) remains insufficiently investigated in the current literature.
+Federated Averaging (FedAvg) (mcmahan2017communication) is perhaps the most widely used FL framework, serving as the foundation for numerous follow-up works that introduce various optimizations.
 Its core principle is to aggregate locally trained client models into a global model, which then initializes the next round of local training.
-Most personalized FL algorithms [45, 30, 35, 8, 7, 10], despite their implementation differences, build upon the FedAvg framework by incorporating personalization mechanisms—and personalized federated LoRA fine-tuning follows the same paradigm.
-For example, the recently proposed FedDPA [45] algorithm personalizes LoRA by introducing both global and local LoRA components.
+Most personalized FL algorithms (yang2024dual; qi2024fdlora; tan2022towards; deng2020adaptive; collins2021exploiting; fallah2020personalized), despite their implementation differences, build upon the FedAvg framework by incorporating personalization mechanisms—and personalized federated LoRA fine-tuning follows the same paradigm.
+For example, the recently proposed FedDPA (yang2024dual) algorithm personalizes LoRA by introducing both global and local LoRA components.
 The global LoRA component is trained using FedAvg, while the local LoRA component is trained independently on each client’s private data.
 This local adaptation can occur either after the global LoRA stabilizes or alternatedly with global LoRA training, in which case the final output is computed as a weighted combination of both components using a pre-defined weighting factor.
-However, our motivational study (see section 2.3) reveals that FedAvg-based approaches may degrade client-specific performance due to harmful cross-client interference.
+However, our motivational study reveals that FedAvg-based approaches may degrade client-specific performance due to harmful cross-client interference.
 We identify two fundamental limitations of FedAvg-based methods like FedDPA in heterogeneous FL settings.
@@
 Distinctly different from FedAvg-based methods, FedALT does not aggregate local models to initialize training rounds.
-Instead, each client continues training on their own previously trained local models, supplemented by a frozen Rest-of-World (RoW) LoRA component.
+Instead, each client continues training on it own previously trained local models, supplemented by a frozen Rest-of-World (RoW) LoRA component.
 The RoW LoRA captures global knowledge without updating during local training rounds, thus avoiding interference with client-specific adaptations.
-To leverage RoW LoRA effectively, we introduce an adaptive mixer, inspired by Mixture-of-Experts (MoE) [18], which dynamically and optimally combines global and local information for each client’s data sample.
+To leverage RoW LoRA effectively, we introduce an adaptive mixer, inspired by Mixture-of-Experts (MoE) (jordan1994hierarchical), which dynamically and optimally combines global and local information for each client’s data sample.
 Our contributions are as follows:
@@
   • We propose FedALT, a novel personalization framework featuring a frozen RoW LoRA and an adaptive mixer, effectively addressing these limitations by dynamically balancing global and local model adaptations.
-  • We perform comprehensive experiments on two widely-used LLMs (Bloom [21] and Llama 2 [38]) using the Flan benchmark [6], demonstrating that FedALT significantly outperforms existing federated LoRA fine-tuning methods.
-2 Preliminaries and Motivations
-2.1 LoRA
+  • We perform comprehensive experiments on two LLMs (Bloom (le2023bloom) and Llama 2 (touvron2023llama2)) using the Flan benchmark (chung2024scaling), demonstrating that FedALT significantly outperforms existing federated LoRA fine-tuning methods.
+Related Work
+Parameter-Efficient Fine-Tuning
+As the size of LLMs continues to grow, full parameter fine-tuning has become increasingly computationally and storage-intensive (hadi2023survey).
+To address this issue, Parameter-Efficient Fine-Tuning (PEFT) methods (ding2023parameter; fu2023effectiveness; han2024parameter; liu2022few; li2021prefix; lester2021power) have been developed to significantly reduce the number of trainable parameters.
+PEFT techniques introduce a limited set of additional trainable parameters to enhance model performance while keeping the majority of pre-trained parameters frozen.
+Among all existing PEFT methods, the most prominent approach is LoRA (hu2021lora), which employs low-rank matrices to approximate the pre-trained weight matrix, updating only the low-rank components.
+LoRA has became a standard method for adapting LLMs like Llama under limited computational resources (lermen2023lora).
+Several works have been proposed to improve LoRA (tian2024hydralora; lermen2023lora; wang2023multilora; sheng2023s; liu2023moelora).
+A recent development, Hydra-LoRA (tian2024hydralora), employs an asymmetric structure with a shared matrix for all samples and distinct matrices for each intrinsic component, thereby enhancing domain adaptation.
+While these advancements primarily focus on centralized learning settings, our work extends the application of PEFT methods to federated LLM fine-tuning, addressing the unique challenges posed by distributed and heterogeneous client data.
+Federated Learning
+The foundational work, FedAvg (mcmahan2017communication), addresses privacy and communication efficiency by aggregating local model to train a shared global model.
+Since then, numerous studies (liuimproving; wang2024aggregation; yang2024fedfed; liu2025fedadamwcommunicationefficientoptimizerconvergence; bian2024accelerating; zhang2025fedel; 10.1145/3746027.3755226; liu2024fedbcgd) have focused on tackling various challenges within FL settings.
+One of the key challenges is data heterogeneity, which makes it difficult to train a single shared global model that performs well across all clients.
+To address this, Personalized Federated Learning (PFL) (tan2022towards; deng2020adaptive; collins2021exploiting; fallah2020personalized) has emerged as an approach to adapt the global model to the specific needs of each client.
+Most PFL research has focused on addressing statistical differences in data distributions (e.g., label distribution skew) and has been applied primarily to smaller, standard models.
+In contrast, our work addresses the unique challenges posed by heterogeneous clients with diverse tasks by leveraging LLM fine-tuning within the FL framework.
+Federated Fine-Tuning
+Several studies (cho2024heterogeneous; wang2024flora; kuang2024federatedscope; wu2024fedbiot) have explored federated fine-tuning approaches.
+(kuang2024federatedscope) proposed federated full parameters fine-tuning, while (sun2022conquering) introduced federated fine-tuning with PEFT using prefix-tuning.
+(zhang2024towards) was the first study to apply LoRA in a federated context.
+(bai2024federated; cho2024heterogeneous) focused on federated fine-tuning where clients have different LoRA ranks, while (wang2024flora) proposed strategies for LoRA initialization to achieve better performance.
+Another studies (Bian_2025_ICCV; sun2024improving) addressed server aggregation bias in LoRA-based federated fine-tuning.
+The most relevant works to ours focus on personalized federated fine-tuning with LoRA (yang2024dual; qi2024fdlora).
+FedDPA (yang2024dual) introduced global and local LoRA components but sufferd from interference during global LoRA training and struggled to balance the two phases effectively.
+PF2LoRA (hao2025personalized) employed a bilevel framework to combine a shared adapter with client-specific adapters but faced interference issues during the aggregation of the shared adapter.
+FDLoRA (qi2024fdlora) utilized personalized LoRA modules to initialize federated training, combining them with a global LoRA module via adaptive fusion; however, it relied on server-side datasets and was prone to interference during global LoRA aggregation, limiting its effectiveness in heterogeneous settings.
+Preliminaries and Motivations
+Low-Rank Adaptation (LoRA)
 Consider a pre-trained model with parameters $\mathbf{W}_{0}\in\mathbb{R}^{l\times d}$, where $\mathbf{W}_{0}$ represents the fixed parameters of the model, and $\mathbf{\Delta W}\in\mathbb{R}^{l\times d}$ denotes the trainable update matrix applied during fine-tuning.
@@
 $\mathbf{W}=\mathbf{W}_{0}+\mathbf{\Delta W}=\mathbf{W}_{0}+\mathbf{B}\mathbf{A},$ (1)
-where $\mathbf{A}$ is typically initialized with random Gaussian values, while $\mathbf{B}$ is initialized to zero, ensuring a stable start to the fine-tuning process.
-2.2 Heterogeneous Federated Fine-Tuning
+where $\mathbf{A}$ is typically initialized with random Gaussian values, while $\mathbf{B}$ is initialized to zero.
+Heterogeneous Federated Fine-Tuning
 Assume there are $K$ institutions (clients), each possessing a distinct local training dataset $\mathcal{D}_{k}=\{X_{k},Y_{k}\}$, where $k$ indexes a client.
@@
 The goal is for each client to learn an individualized fine-tuned model $\textbf{W}_{k}$ that best fits its local dataset.
-The objective of the training process can be formulated as:
+The objective can be formulated as:
 $\min_{\mathcal{W}}\frac{1}{K}\sum_{k=1}^{K}L_{k}(X_{k},Y_{k},\mathbf{W}_{k}),$ (2)
-where $L_{k}(\cdot)$ represents the loss function for client $k$, and $\mathcal{W}$ denotes the set of personalized fine-tuned models $\{\mathbf{W}_{k}\}_{k=1}^{K}$.
-2.3 Motivational Study
+where $L_{k}(\cdot)$ represents the loss function for client $k$, and $\mathcal{W}$ denotes the set of personalized models $\{\mathbf{W}_{k}\}_{k=1}^{K}$.
+Motivational Study
 Federated fine-tuning operates in highly heterogeneous environments where client tasks can differ significantly.
@@
 Fact 1: FL Can Transfer Useful Knowledge but Also Introduces Harmful Interference.
-In FL settings, the data collected by each client can be highly diverse [17, 28, 40].
-We conducted a motivational experiment using the Flan dataset [6], selecting 8 tasks and creating 8 clients, each specializing in a single task.
+In FL settings, the data collected by each client can be highly diverse (huang2022learn; mendieta2022local; wang2024taming).
+We conducted a motivational experiment using the Flan dataset (chung2024scaling), selecting 8 tasks and creating 8 clients, each specializing in a single task.
 To demonstrate potential interference in federated fine-tuning, we compared two baseline methods.
-The first method, FedIT [49], combines FedAvg with LoRA fine-tuning.
+The first method, FedIT (zhang2024towards), combines FedAvg with LoRA fine-tuning.
 The second method is Local Only LoRA fine-tuning, where each client fine-tunes its LoRA locally without any communication or information sharing between clients.
-Figure 2: Motivational study results. See details in section 2.3.
 The results, shown in fig.
@@
 This dual outcome underscores both the promise and the pitfalls of federated fine-tuning in heterogeneous settings.
+Figure 2: Motivational study results.
 Fact 2: FedAvg Paradigm Struggles to Mitigate Interference
 Having established the existence of harmful interference, we next explored potential mitigation strategies.
-In centralized LoRA fine-tuning, HydraLoRA [36] found that splitting a single large LoRA into multiple smaller LoRAs reduces cross-domain interference.
+In centralized LoRA fine-tuning, HydraLoRA (tian2024hydralora) found that splitting a single large LoRA into multiple smaller LoRAs reduces cross-domain interference.
 However, FL introduces additional challenges, as data domains are distributed across clients.
@@
 Consequently, direct adaptations of centralized interference-mitigation strategies fail to address the core issue in FL.
-3 Proposed Method: FedALT
+Proposed Method: FedALT
 Our motivational studies reveal that harmful interference in FL arises from server aggregation and the use of the aggregated model to initialize subsequent training rounds.
@@
 This approach eliminates interference entirely but fails to leverage shared knowledge from other clients.
-To address this trade-off, we propose a novel personalized federated LoRA fine-tuning algorithm that fundamentally diverges from the FedAvg paradigm, prioritizing local training while selectively integrating useful global information.
+To address this trade-off, we propose a novel personalized federated LoRA fine-tuning algorithm that diverges from the FedAvg paradigm, prioritizing local training while selectively integrating useful global information.
 Our method introduces two key innovations: 1.
@@
 In the following of this section, we detail the design and functionality of these components.
-Additionally, a theoretical convergence analysis of the proposed method is provided in the appendix F.
-3.1 Individual and Rest-of-World (RoW) LoRA
-Each client $k$ maintains two types of LoRAs: an Individual LoRA, denoted by $\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$, and a Rest-of-World (RoW) LoRA, denoted by $\mathbf{A}^{R}_{k}/\mathbf{B}^{R}_{k}$.
+The theoretical convergence analysis is provided in the supplementary material.
+Individual and Rest-of-World (RoW) LoRA
+Each client $k$ maintains two types of LoRAs: an Individual LoRA, denoted by $\mathbf{A}^{L}_{k}/\mathbf{B}^{L}_{k}$, and a RoW LoRA, denoted by $\mathbf{A}^{R}_{k}/\mathbf{B}^{R}_{k}$.
 The Individual LoRA captures information obtained locally by the client and is isolated from the information of other clients.
@@
 Each client then determines its RoW LoRA by subtracting its individual LoRA from the global average LoRA.
-Now, consider the scenario where a weight $\frac{1}{K}$ is assigned to the local LoRA and $1-\frac{1}{K}$ to the RoW LoRA when combining their outputs.
-In this case, the combined LoRA is mathematically equivalent to the averaged LoRA.
-This equivalence arises naturally from the definitions of the RoW and averaged LoRAs.
-If these two LoRAs are trained simultaneously on local datasets, the approach essentially reduces to the standard FedAvg (or FedIT) algorithm, as all clients would end up with the same combined LoRA.
-However, it is crucial to emphasize that the RoW LoRA is not further updated on clients using their local datasets.
+It is crucial to emphasize that the RoW LoRA is not further updated on clients using their local datasets.
 Instead, it is only updated through averaging at the end of each learning round.
@@
 Furthermore, skipping local training for the RoW LoRA reduces the computational workload on clients by half.
-3.2 Dynamic Mixture-of-Experts Mechanism
+Dynamic Mixture-of-Experts Mechanism
 A natural question arises: why not simply add the RoW LoRA directly to the pre-trained model, i.e., $\mathbf{W}_{0}\leftarrow\mathbf{W}_{0}+\mathbf{B}^{R}_{k}\mathbf{A}^{R}_{k}$, since the RoW LoRA remains fixed during local training?
@@
 Given the complexity of modern LLMs, where LoRA modules are inserted at multiple layers, learning $\alpha(x)$ is non-trivial.
-Inspired by [18], we adopt a Mixture-of-Experts (MoE) approach to compute $\alpha(x)$.
+Inspired by (jordan1994hierarchical), we adopt a Mixture-of-Experts (MoE) approach to compute $\alpha(x)$.
 Specifically, we introduce a mixer $\mathbf{G}_{k}\in\mathbb{R}^{2\times d}$ which is a dense layer with trainable weights (a transformation matrix), followed by a softmax function, to learn the weighting:
@@
 This personalization ensures that the weight adjustments reflect each client’s unique data distribution and learning objectives.
-3.3 Algorithm Workflow
+Algorithm Workflow
 By combining the two LoRA modules (Individual LoRA and RoW LoRA) with dynamic weights, our proposed method achieves a balance between leveraging useful information from other clients and isolating harmful interference.
-To better illustrate the FedALT training process, we provide an overview of the server and client operations during fine-tuning at round $t$.
+To better illustrate the FedALT training process, we provide an overview of the server and client operations at round $t$.
 Server Side.
@@
 Once the local fine-tuning is complete, the updated Individual LoRA is uploaded to the server, while $\mathbf{G}_{k}$ stays local to the client.
-The theoretical analysis of the convergence of the proposed FedALT is presented in appendix F.
-3.4 Discussion: Why a Single RoW LoRA?
-FedALT prioritizes local training while selectively incorporating useful information from other clients via a dynamic weighting mechanism.
-Ideally, each client could tailor this incorporation by maintaining separate LoRA components for every other client.
-In this approach, client $k$ would receive the local LoRA components $\mathbf{A}^{L}_{j}/\mathbf{B}^{L}_{j}$ from all other clients $j\neq k$, freeze them during local training, and learn separate weights for each $j$:
-$y=\mathbf{W}_{0}x+\alpha_{k,k}(x)\mathbf{B}^{L}_{k}\mathbf{A}^{L}_{k}x+\sum_{j\neq k}\alpha_{k,j}(x)\mathbf{B}^{L}_{j}\mathbf{A}^{L}_{j}x$
-where $\sum_{j}\alpha_{k,j}(x)=1$.
-This would allow client $k$ to capture the distinct influence of each other client’s LoRA.
-However, this approach presents significant drawbacks.
-First, each client would need to store and communicate a LoRA component for every other client, leading to memory and bandwidth overhead that scales linearly with the number of clients $K$, making it impractical for large-scale settings.
-Second, training the mixer becomes increasingly complex with more LoRA components, requiring additional parameters and longer convergence times.
-To address these challenges, we adopt a single RoW LoRA, which aggregates knowledge from other clients into a single component.
-While this reduces computational and communication costs, it may limit personalization.
-Future work could explore clustering-based approaches to maintain a small number of RoW LoRA components, capturing finer-grained client relationships while remaining scalable.
-4 Experiments
-In this section, we evaluate and compare the performance of our proposed method, FedALT, with other baseline methods. Detailed training implementations are provided in the appendix B.
-Datasets.
-Our training datasets are derived from Flan [6], a comprehensive benchmark comprising over 60 NLP datasets across more than 12 diverse task types.
-To simulate client heterogeneity in the main experiments, we assume 8 clients, each assigned a distinct NLP task dataset.
-Sensitivity analysis with varying numbers of clients and degrees of task heterogeneity are presented in section 4.3.
-Detailed dataset descriptions, including task distributions and client assignments, are provided in appendix A.
-Pre-trained Model and Baselines.
-We utilize LLaMA2-7B [38] and Bloom-560M [21] as the pre-trained models for all baseline and proposed fine-tuning methods.
-To evaluate the performance of our proposed method, we compare it with the state-of-the-art methods in federated fine-tuning with LoRA, including FedIT [49], FFA-LoRA [34].
-Additionally, we compare it with state-of-the-art methods focused on personalized federated fine-tuning with LoRA, such as FedDPA [45], PF2LoRA [15], FedSA [12], and FDLoRA [30].
-We also include the Local Only baseline, where clients fine-tune their models independently without federated collaboration.
-Detailed descriptions of these baseline methods are provided in the appendix C.
-Table 1: Performance comparison with baseline methods with LLaMA2-7B. The average column represents the overall performance across these tasks. See details in section 4.1.
-Methods LLaMA2-7B
-Commonsense Coreference Natural Language Paraphrase Reading Sentiment Struct to Text Average
+Methods Commonsense Coreference Natural Language Paraphrase Reading Sentiment Struct to Text Average
 Reasoning Resolution Inference Detection Comprehension Analysis Text Classification
+LLaMA2-7B
 Local 73.83 74.62 58.73 71.00 55.82 46.56 55.14 67.18 62.86
@@
 FDLoRA 76.29 75.60 66.12 75.67 57.39 49.85 52.85 67.59 65.17
-FedALT 76.12 83.04 67.73 77.80 59.41 51.57 53.12 71.60 67.55
-Table 2: Performance comparison across various NLP tasks with Bloom-560M. The average column represents the overall performance across these tasks. See details in section 4.1.
-Methods Bloom-560M
-Commonsense Coreference Natural Language Paraphrase Reading Sentiment Struct to Text Average
-Reasoning Resolution Inference Detection Comprehension Analysis Text Classification
+[…]
+Bloom-560M
 Local 52.13 39.20 37.55 62.67 29.45 42.91 41.33 49.99 44.40
@@
 FDLoRA 53.65 40.22 37.27 68.89 29.95 42.89 35.39 56.05 45.54
-FedALT 56.39 41.27 44.45 70.67 30.62 43.04 39.24 59.03 48.09
-4.1 Performance Comparisons
+[…]
+Table 1: Performance comparison with baseline methods across different models.
+Experiments
+Datasets.
+Our training datasets are derived from Flan (chung2024scaling), a comprehensive benchmark comprising over 60 NLP datasets across more than 12 diverse task types.
+To simulate client heterogeneity in the main experiments, we assume 8 clients, each assigned a distinct NLP task dataset.
+Detailed descriptions, including task distributions and client assignments, are provided in supplementary material.
+Pre-trained Model and Baselines.
+We utilize LLaMA2-7B (touvron2023llama2) and Bloom-560M (le2023bloom) as the pre-trained models for all baseline and proposed fine-tuning methods.
+To evaluate the performance of our proposed method, we compare it with the state-of-the-art methods in federated fine-tuning with LoRA, including FedIT (zhang2024towards), FFA-LoRA (sun2024improving).
+Additionally, we compare it with state-of-the-art methods focused on personalized federated fine-tuning with LoRA, such as FedDPA (yang2024dual), PF2LoRA (hao2025personalized), FedSA (guo2024selective), and FDLoRA (qi2024fdlora).
+We also include the Local Only baseline, where clients fine-tune their models independently without federated collaboration.
+Detailed descriptions of these baseline methods are provided in the supplementary material.
+(a) Impact of RoW LoRA. Isolating client-specific information is more effective than using globally averaged updates.
+(b) Impact of dynamic weighting via MOE. The personalized mixer outperforms fixed or shared weighting.
+Figure 3: Ablation studies of FedALT.
+(a) Effect of number of clients.
+(b) Effect of LoRA rank.
+(c) Effect of local epochs.
+Figure 4: Sensitivity Analyses of FedALT under different training configurations.
+Performance Comparisons
 We compare the performance of FedALT, with other baseline methods based on each client’s performance on its specific test dataset, aligned with its training dataset.
@@
 The consistent performance gains observed with both LLaMA2-7B and Bloom-560M confirm the generalizability of our approach across diverse model families and parameter scales.
-4.2 Ablation Study
-Table 3: Impact of decoupling LoRA training. See Sec. 4.2.
+Ablation Study
 Methods Same Trainable Same Inference Avg. Perf. RoW-Update (rank = 4) $\checkmark$ $\times$ 63.59 RoW-Update (rank = 8) $\times$ $\checkmark$ 65.75 FedALT 67.55
+Table 2: Impact of decoupling LoRA training.
 Impact of Decoupling LoRA Training.
@@
 The key distinction lies in whether the RoW LoRA is updated locally.
-As shown in table 3, while the second method outperforms the first due to its increased capacity, both alternatives underperform relative to FedALT.
+As shown in table 2, while the second method outperforms the first due to its increased capacity, both alternatives underperform to FedALT.
 Notably, even with more trainable parameters, the second method fails to match FedALT’s performance, highlighting the benefits of decoupling local updates from the global component.
@@
 In the second fixed strategy, we set the contribution weight to $\alpha=1/K$, where $K$ is the number of clients.
-This configuration effectively reduces the method to using the averaged LoRA, relying solely on each client’s last-round LoRA modules for the next round’s initialization.
 The results in fig.
@@
 The results confirm that retaining a personalized $\mathbf{G}_{k}$ for each client yields optimal performance, emphasizing the importance of input-specific and client-specific adjustments.
-(a) Impact of RoW LoRA. Isolating client-specific information is more effective than using globally averaged updates. See section 4.2.
-(b) Impact of dynamic weighting via MOE. The personalized mixer outperforms fixed or shared weighting. See section 4.2.
-Figure 3: Ablation studies of FedALT.
-4.3 Sensitivity Analysis
-Effect of Data Heterogeneity.
-To further evaluate the robustness of our method under varying degrees of client heterogeneity, we design experiments across three settings.
-Our main experimental configuration, where each client is assigned a unique task, is denoted as the High Heterogeneous setting.
-We also include two alternative configurations.
-In the Low Heterogeneous setting, 8 clients are divided between only two tasks—Sentiment Analysis and Natural Language Inference—with clients sharing a task drawing data from the same source, resulting in minimal heterogeneity.
-The Mild Heterogeneous setting involves 8 clients assigned datasets spanning a total of 6 distinct NLP tasks.
-As shown in table 4, FedALT consistently outperforms all baseline methods across all levels of heterogeneity, using both Bloom and LLaMA2 as pre-trained models.
-These results highlight the robustness and adaptability of FedALT in handling diverse and heterogeneous data distributions.
-Due to space limitations, we report only the average results here; detailed per-client performance can be found in the appendix E.
+Sensitivity Analysis
 Effect of Number of Clients.
@@
 Additionally, the server-side computation of the RoW LoRA remains lightweight, as it involves only simple averaging operations.
-Table 4: Performance under varying heterogeneity levels. Evaluation on Bloom-560M and LLaMA2-7B models across three levels of task heterogeneity. See section 4.3.
-Methods Bloom-560M LLaMA2-7B Low Mild High Low Mild High Local 48.65 50.21 44.40 60.59 66.67 62.86 FedIT 52.15 49.03 44.33 62.24 64.53 62.19 FFA-LoRA 51.33 47.32 40.44 61.07 62.15 56.78 FedSA 53.03 52.55 46.98 62.87 67.44 63.47 FedDPA 53.12 52.06 46.99 62.06 67.49 64.64 PF2LoRA 51.10 51.97 45.87 63.56 67.60 64.44 FDLoRA 52.51 51.42 45.54 62.42 66.56 65.17 FedALT 54.48 54.89 48.09 64.66 70.48 67.55
 Effect of LoRA Rank.
@@
 Effect of Local Epochs.
-We also investigate the effect of varying local training epochs (5 and 10), while keeping all other configurations unchanged.
+We investigate the effect of varying local training epochs (5 and 10), while keeping all other configurations unchanged.
 As presented in fig.
-4(c), our method outperforms all baselines under both settings, further highlighting its robustness to different local training epochs.
-(a) Effect of number of clients. See section 4.3.
-(b) Effect of LoRA rank. See section 4.3.
-(c) Effect of local epochs. See section 4.3.
-Figure 4: Sensitivity Analyses of FedALT under different training configurations.
-5 Related Work
-FedAvg [27] addresses privacy and communication efficiency in federated learning by aggregating locally trained model parameters to construct a shared global model.
-With the growing interest in LLMs, FL has emerged as a promising framework for fine-tuning LLMs while preserving data privacy.
-Recent works have explored federated fine-tuning of LLMs by either updating the entire model [20] or applying PEFT techniques such as prefix-tuning [33] and LoRA [49].
-Subsequent studies [1, 5, 43, 3, 34] have tackled key challenges in this domain, including heterogeneous LoRA configurations, initialization strategies, and server-side aggregation bias.
-Personalized federated fine-tuning methods [15, 45, 30] have proposed combining global and local LoRA modules, but often suffer from cross-client interference during aggregation or depend on server-side data.
-Our method introduces a robust and scalable approach to personalized fine-tuning across heterogeneous clients.
-A more comprehensive review of related work is provided in appendix C.
-6 Conclusion
+4(c), our method outperforms all baselines, further highlighting its robustness to different local training epochs.
+Additional experiments on the effect of data heterogeneity can be found in the supplementary material.
+Conclusion
 In this work, we introduced FedALT, a personalized federated LoRA fine-tuning framework that departs from conventional FedAvg paradigms.
@@
 Future research directions include investigating clustering-based approaches for maintaining multiple RoW LoRA components, and extending our framework to multi-modal foundation models.
+Acknowledgments
+The work of Jieming Bian, Lei Wang and Jie Xu is partially supported by NSF under grants 2433886, 2505381 and 2515982.
+The work of Letian Zhang is partially supported by NSF under grant 2348279 and also supported by MTSU Stark Land project.
 Appendix A Datasets
-In this paper, we follow the setup proposed in [45] to construct federated datasets for heterogeneous client settings using the Flan collection [6].
-The Flan benchmarks comprises over 12 NLP task categories, each containing multiple datasets.
+In this paper, we follow the setup proposed in (yang2024dual) to construct federated datasets for heterogeneous client settings using the Flan collection (chung2024scaling).
+The Flan benchmarks comprise over 12 NLP task categories, each containing multiple datasets.
 These tasks, derived from diverse contextual domains, inherently involve complex distribution shifts.
@@
 Specifically, in the 16-client configuration, each client receives 300 training samples and 150 testing samples, while in the 24-client setting, each client is allocated 200 training samples and 100 testing samples.
-The specific tasks and datasets used in each configuration are provided in tables 5, 6 and 7.
-While some datasets appear in multiple configurations, the sample selections are made independently and randomly.
+The specific tasks and datasets used in each configuration are provided in Tables 3, 4, 5.
+Although some datasets appear in multiple configurations, sample selections are made independently and randomly.
 The original Flan tasks exhibit highly diverse formats, making standardized preprocessing and evaluation across LLMs challenging.
-Even within a single dataset (e.g., Sentiment140), sample formats can vary significantly, as illustrated in tables 8 and 9.
-To ensure consistency, we adopt a unified prompt template based on [45], shown in table 10.
-For evaluation, we follow FedDPA [45] and use ROUGE-1 as the metric across all tasks, providing a unified measure of model performance.
+Even within a single dataset (e.g., Sentiment140), sample formats can vary significantly, as illustrated in tables 6 and 7.
+To ensure consistency, we adopt a unified prompt template based on (yang2024dual), shown in table 8.
+For evaluation, we follow FedDPA (yang2024dual) and use ROUGE-1 as the metric across all tasks, providing a unified measure of model performance.
 This choice is particularly appropriate as Flan inherently structures NLP tasks in a generative format, where ROUGE-1 effectively captures lexical overlap between generated outputs and reference answers.
@@
 While task-specific metrics might offer more nuanced evaluation for individual tasks, they would introduce incomparability issues that complicate assessment of the overall federated learning performance, making ROUGE-1 a methodologically sound choice.
-Table 5: Tasks and Datasets for High Heterogeneous
 Task Dataset Description Text Classification ag_news_subset News classification into categories.
@@
 Coreference Resolution definite_pronoun_resolution Resolve ambiguous pronouns.
-Table 6: Tasks and Datasets for Mild Heterogeneous
+Table 3: Tasks and Datasets for High Heterogeneous
 Task Dataset Description Struct to Text 1 common_gen Generate text from concepts.
@@
 Sentiment Analysis sst2 Analyze sentiment of text.
-Table 7: Tasks and Datasets for Low Heterogeneous
+Table 4: Tasks and Datasets for Mild Heterogeneous
 Task Dataset Description Sentiment Analysis 1 sst2 Analyze sentiment of short movie reviews.
@@
 Natural Language Inference 2 snli Determine relationships between sentence pairs.
-Table 8: Examples of Data in Sentiment Analysis Task.
+Table 5: Tasks and Datasets for Low Heterogeneous
 Data Examples Instruction: @celinieee The crafts thingy is already done.
@@
 Dataset: sentiment140
-Table 9: Examples of Data in Struct to Text Task.
+Table 6: Examples of Data in Sentiment Analysis Task.
 Data Examples Instruction: What are the most important words in the following sentence: pig in a pen on a pig farm Output: farm, pen, pig Dataset: common_gen Instruction: Generate a sentence about this data: Wildwood eatType coffee shop; Wildwood priceRange moderate; Wildwood customer rating 1 out of 5; Wildwood near Ranch Output: The Wildwood coffee shop is an average-priced location that serves a light menu of food and alcohol.
@@
 Dataset: dart
+Table 7: Examples of Data in Struct to Text Task.
+Template
+Prompt Input
+Instruction: {instruction}
+Response:
 Table 8: Prompt Template.
-Template
-Prompt Input Instruction: {instruction} Response:
 Appendix B Training Implementations
@@
 For the Local Only baseline, the total number of training rounds is set to 50.
-The learning rate is explored in the range $\{3\times 10^{-2},1\times 10^{-3},3\times 10^{-3},1\times 10^{-4},3\times 10^{-4},1\times 10^{-5}\}$, and the AdamW optimizer is utilized, following [45].
+The learning rate is explored in the range $\{3\times 10^{-2},1\times 10^{-3},3\times 10^{-3},1\times 10^{-4},3\times 10^{-4},1\times 10^{-5}\}$, and the AdamW optimizer is utilized, following (yang2024dual).
 LoRA modules are inserted into the $q$ and $v$ components of each self-attention layer.
@@
 Due to computational constraints, the pre-trained LLaMA2 model is loaded using 8-bit quantization to reduce memory requirements while maintaining performance.
-Appendix C Related Work and Compared Methods
-C.1 Parameter-Efficient Fine-Tuning
-As the size of LLMs continues to grow, full parameter fine-tuning has become increasingly computationally and storage-intensive [13].
-To address this issue, Parameter-Efficient Fine-Tuning (PEFT) methods [9, 11, 14, 25, 24, 23] have been developed to significantly reduce the number of trainable parameters.
-PEFT techniques introduce a limited set of additional trainable parameters to enhance model performance while keeping the majority of pre-trained parameters frozen.
-Among all existing PEFT methods, the most prominent approach is LoRA [16], which employs low-rank matrices to approximate the pre-trained weight matrix, updating only the low-rank components.
-LoRA has gained significant attention recently, becoming a standard method for adapting LLMs like Llama under limited computational resources [22].
-Several works have been proposed to improve LoRA [36, 22, 41, 32, 26].
-For instance, a recent development, Hydra-LoRA [36], employs an asymmetric structure with a shared matrix for all samples and distinct matrices for each intrinsic component, thereby enhancing domain adaptation.
-While these advancements primarily focus on centralized learning settings, our work extends the application of PEFT methods to federated LLM fine-tuning, addressing the unique challenges posed by distributed and heterogeneous client data.
-C.2 Federated Learning
-The foundational work in federated learning, FedAvg [27], addresses privacy and communication efficiency by aggregating local model parameters to train a shared global model.
-Since then, numerous studies [39, 42, 46, 2] have focused on tackling various challenges within FL settings.
-One of the key challenges is data heterogeneity, which makes it difficult to train a single shared global model that performs well across all clients.
-To address this, Personalized Federated Learning(PFL) [35, 8, 7, 10] has emerged as an approach to adapt the global model to the specific needs of each client in the federation.
-Most PFL research has focused on addressing statistical differences in data distributions (e.g., label distribution skew) and has been applied primarily to smaller, standard models.
-In contrast, our work addresses the unique challenges posed by heterogeneous clients with diverse NLP tasks by leveraging LLM fine-tuning within the FL framework.
-C.3 Federated Fine-Tuning
-Several studies [1, 5, 43, 20, 44, 3] have explored federated fine-tuning approaches.
-[20] proposed federated fine-tuning with all parameters updated, while [33] introduced federated fine-tuning with PEFT using prefix-tuning.
-[49] was the first study to apply LoRA in a federated context.
-[1, 5] focused on federated fine-tuning where clients have different LoRA ranks, while [43] proposed strategies for LoRA initialization to achieve better performance.
-Another studies [3, 34] addressed server aggregation bias in LoRA-based federated fine-tuning.
-The most relevant works to ours focus on personalized federated fine-tuning with LoRA [15, 45, 30].
-FedDPA [45] introduced global and local LoRA components but sufferd from interference during global LoRA training and struggled to balance the two phases effectively.
-PF2LoRA [15] employed a bilevel framework to combine a shared adapter with client-specific adapters but faced interference issues during the aggregation of the shared adapter.
-FDLoRA [30] utilized personalized LoRA modules to initialize federated training, combining them with a global LoRA module via adaptive fusion; however, it relied on server-side datasets and was prone to interference during global LoRA aggregation, limiting its effectiveness in heterogeneous settings.
+Appendix C Compared Methods
 A detailed overview of the baseline methods used for comparison in our experiments is provided below.
-FedIT [49].
+FedIT (zhang2024towards).
 FedIT integrates LoRA with the FedAvg framework, making it one of the earliest approaches to combine these techniques.
@@
 This interference often reduces its efficiency and, in some cases, results in worse performance than the Local Only method.
-FFA-LoRA [34].
+FFA-LoRA (sun2024improving).
 FFA-LoRA aims to reduce communication costs by fixing the randomly initialized non-zero $\mathbf{A}$ matrices and fine-tuning only the zero-initialized $\mathbf{B}$ matrices.
 While this approach lowers communication overhead, it impairs the learning capacity of LoRA, as fixed $\mathbf{A}$ matrices limit the model’s ability to adapt to diverse client data, resulting in suboptimal performance.
-FedSA [12].
+FedSA (guo2024selective).
 FedSA introduces an asymmetric approach to LoRA matrix updates in federated learning, where only the $\mathbf{A}$ matrices are shared with the server for aggregation, while the $\mathbf{B}$ matrices remain local to capture client-specific knowledge.
@@
 The failure to leverage this knowledge effectively results in reduced overall performance.
-PF2LoRA [15].
+PF2LoRA (hao2025personalized).
 PF2LoRA adopts a two-level low-rank adaptation framework, where a common adapter shared across all clients is combined with client-specific lightweight adapters for personalization.
@@
 However, the aggregation of the shared adapter during the server aggregation step can reintroduce harmful interference, limiting the method’s effectiveness in highly heterogeneous settings.
-FDLoRA [30].
+FDLoRA (qi2024fdlora).
 FDLoRA follows a two-step process where personalized LoRA modules are first trained locally and then used to initialize federated training to obtain a global LoRA module.
@@
 Moreover, the aggregation step during global LoRA training remains prone to interference, reducing its robustness in handling client heterogeneity.
-FedDPA [45].
+FedDPA (yang2024dual).
 FedDPA is the most closely related to our proposed method, as it also targets settings where clients possess heterogeneous tasks (not just label non-IID data).
@@
 In contrast, FedALT removes the need for phase separation and additional hyperparameters, simplifying training and improving robustness.
-Appendix D The Algorithm of FedALT
+Appendix D Discussion: Why a Single RoW LoRA?
+FedALT prioritizes local training while selectively incorporating useful information from other clients via a dynamic weighting mechanism.
+Ideally, each client could tailor this incorporation by maintaining separate LoRA components for every other client.
+In this approach, client $k$ would receive the local LoRA components $\mathbf{A}^{L}_{j}/\mathbf{B}^{L}_{j}$ from all other clients $j\neq k$, freeze them during local training, and learn separate weights for each $j$:
+$y=\mathbf{W}_{0}x+\alpha_{k,k}(x)\mathbf{B}^{L}_{k}\mathbf{A}^{L}_{k}x+\sum_{j\neq k}\alpha_{k,j}(x)\mathbf{B}^{L}_{j}\mathbf{A}^{L}_{j}x$
+where $\sum_{j}\alpha_{k,j}(x)=1$.
+This would allow client $k$ to capture the distinct influence of each other client’s LoRA.
+However, this approach presents significant drawbacks.
+First, each client would need to store and communicate a LoRA component for every other client, leading to memory and bandwidth overhead that scales linearly with the number of clients $K$, making it impractical for large-scale settings.
+Second, training the mixer becomes increasingly complex with more LoRA components, requiring additional parameters and longer convergence times.
+To address these challenges, we adopt a single RoW LoRA, which aggregates knowledge from other clients into a single component.
+While this reduces computational and communication costs, it may limit personalization.
+Future work could explore clustering-based approaches to maintain a small number of RoW LoRA components, capturing finer-grained client relationships while remaining scalable.
+Appendix E The Algorithm of FedALT
 We summarize the pseudocode of FedALT in Algorithms 1.
-Input: Pretrained model $W_{0}$, number of clients $K$, communication rounds $T$
-1 for each client $k$ do
-2       Initialize Individual LoRA $(A_{k}^{L},B_{k}^{L})$ and Mixer $G_{k}$ randomly;
-3       Initialize RoW LoRA $(A_{k}^{R},B_{k}^{R})\leftarrow 0$;
-4
-5for each round $t=1$ to $T$ do
-       // Client-side training
-6       for each client $k$ in parallel do
-7             Receive RoW LoRA $(A_{k}^{R},B_{k}^{R})$ from server;
-8             for each local epoch do
-9                   Update $(A_{k}^{L},B_{k}^{L})$ and Mixer $G_{k}$ using gradient descent;
-                   // Model output:
-10                   $y=W_{0}x+\alpha_{k}(x)B_{k}^{L}A_{k}^{L}x+(1-\alpha_{k}(x))B_{k}^{R}A_{k}^{R}x$;
-11                   where $\alpha_{k}(x)=\text{softmax}(G_{k}x)$
-12            Upload $(A_{k}^{L},B_{k}^{L})$ to server;
-13
-14
-      // Server-side aggregation
-15       for each client $k$ do
-16             Compute RoW LoRA: $A_{k}^{R}\leftarrow\frac{1}{K-1}\sum_{m\neq k}A_{m}^{L}$, $B_{k}^{R}\leftarrow\frac{1}{K-1}\sum_{m\neq k}B_{m}^{L}$;
-17             Send $(A_{k}^{R},B_{k}^{R})$ to client $k$;
-18
-19
 Algorithm 1 FedALT
-Appendix E Additional Experiment Results
-E.1 Training Workload
+0: Pretrained model $W_{0}$, number of clients $K$, communication rounds $T$
+1:for each client $k$ do
+2:  Initialize LoRA $(A_{k}^{L},B_{k}^{L})$ and mixer $G_{k}$ randomly
+3:  Initialize RoW LoRA $(A_{k}^{R},B_{k}^{R})\leftarrow 0$
+4:end for
+5:for $t=1$ to $T$ do
+6:  Client-side training:
+7:  for each client $k$ in parallel do
+8:   Receive RoW LoRA $(A_{k}^{R},B_{k}^{R})$ from server
+9:   for each local epoch do
+10:    Update $(A_{k}^{L},B_{k}^{L})$ and $G_{k}$ using gradient descent
+11:    Compute output:
+12:    $y=W_{0}x+\alpha_{k}(x)B_{k}^{L}A_{k}^{L}x+(1-\alpha_{k}(x))B_{k}^{R}A_{k}^{R}x$
+13:    where $\alpha_{k}(x)=\text{softmax}(G_{k}x)$
+14:   end for
+15:   Upload $(A_{k}^{L},B_{k}^{L})$ to server
+16:  end for
+17:  Server-side aggregation:
+18:  for each client $k$ do
+19:   $A_{k}^{R}\leftarrow\frac{1}{K-1}\sum_{m\neq k}A_{m}^{L}$
+20:   $B_{k}^{R}\leftarrow\frac{1}{K-1}\sum_{m\neq k}B_{m}^{L}$
+21:   Send $(A_{k}^{R},B_{k}^{R})$ to client $k$
+22:  end for
+23:end for
+Appendix F Additional Experiment Results
+Training Workload
 In table 9, we compare the training workload of our proposed FedALT with existing baselines using the Bloom-560M model.
@@
 These comparisons validate two core design choices in FedALT: the effectiveness of decoupled training and the benefit of dynamic mixing for personalized adaptation.
+Methods Trainable Parameters Avg Local Epoch Time Inference Time FedIT 0.0622% 6.66 s 8.89 s FFA-LoRA 0.0311% 6.56 s 8.77 s FedALT 0.0661% 7.05 s 8.94 s
 Table 9: Comparison of training workload. FedALT uses a comparable number of LoRA parameters to other baselines.
-Methods Trainable Parameters Avg Local Epoch Time Inference Time FedIT 0.0622% 6.66 s 8.89 s FFA-LoRA 0.0311% 6.56 s 8.77 s FedALT 0.0661% 7.05 s 8.94 s
+Methods Trainable Parameters Avg Performance RoW-Update (rank=8) 0.1283% 65.75 Fixed Weight (rank=16) 0.1244% 64.59 FedALT 0.0661% 67.55
 Table 10: Comparison of trainable parameters and performance. FedALT achieves the highest performance while using fewer parameters.
-Methods Trainable Parameters Avg Performance
-RoW-Update (rank=8) 0.1283% 65.75
-Fixed Weight (rank=16) 0.1244% 64.59
-FedALT 0.0661% 67.55
-E.2 Per-Client Performance
-Due to space limitations, we reported only the average results for the Low Heterogeneous and Mild Heterogeneous settings in the main paper.
-Here, we present the detailed per-client performance in tables 13, 14, 15 and 16.
-Table 13: Performance on mild heterogeneous (LLaMA2-7B). Evaluation across tasks in Table 6.
+Effect of Data Heterogeneity
+To further evaluate the robustness of our method under varying degrees of client heterogeneity, we design experiments across three settings.
+Our main experimental configuration, where each client is assigned a unique task, is denoted as the High Heterogeneous setting.
+We also include two alternative configurations.
+In the Low Heterogeneous setting, 8 clients are divided between only two tasks—Sentiment Analysis and Natural Language Inference—with clients sharing a task drawing data from the same source, resulting in minimal heterogeneity.
+The Mild Heterogeneous setting involves 8 clients assigned datasets spanning a total of 6 distinct NLP tasks.
+As shown in tables 11, 12, 13 and 14., FedALT consistently outperforms all baseline methods across all levels of heterogeneity, using both Bloom and LLaMA2 as pre-trained models.
+These results highlight the robustness and adaptability of FedALT in handling diverse and heterogeneous data distributions.
 Methods LLaMA2-7B Average
@@
 FDLoRA 52.71 58.37 74.74 69.62 66.27 71.00 65.99 73.80 66.56
-FedALT 51.73 60.72 81.91 74.51 71.12 80.33 70.28 73.22 70.48
-Table 14: Performance on mild heterogeneous (Bloom-560M). Evaluation across tasks in Table 6.
+[…]
+Table 11: Performance on mild heterogeneous (LLaMA2-7B). Evaluation across tasks in Table 4.
 Methods Bloom-560M Average
@@
 FDLoRA 40.66 44.62 38.91 43.54 51.17 70.33 54.13 68.00 51.42
-FedALT 41.10 48.64 39.80 48.14 56.63 71.27 64.84 68.67 54.89
-Table 15: Performance on low heterogeneous (LLaMA2-7B). Evaluation across 8 clients split by task-dataset pairs. SA1–SA2 use sst2, SA3–SA4 use sentiment140, NLI1–NLI2 use qnli, and NLI3–NLI4 use snli.
+[…]
+Table 12: Performance on mild heterogeneous (Bloom-560M). Evaluation across tasks in Table 4.
 Methods SA1 SA2 SA3 SA4 NLI1 NLI2 NLI3 NLI4 Average
@@
 FDLoRA 68.54 65.46 65.93 63.36 47.09 49.19 70.69 69.07 62.42
-FedALT 73.05 67.86 65.91 72.82 47.29 50.01 70.42 69.88 64.66
-Table 16: Performance on low heterogeneous (Bloom-560M). Each dataset is split across two clients. See Table 7 for dataset mapping.
+[…]
+Table 13: Performance on low heterogeneous (LLaMA2-7B). Evaluation across 8 clients split by task-dataset pairs. SA1–SA2 use sst2, SA3–SA4 use sentiment140, NLI1–NLI2 use qnli, and NLI3–NLI4 use snli.
 Methods SA1 SA2 SA3 SA4 NLI1 NLI2 NLI3 NLI4 Average
@@
 FDLoRA 50.66 46.72 49.38 54.07 42.35 45.82 64.63 66.45 52.51
-FedALT 54.06 49.18 53.34 55.78 42.94 45.51 67.57 67.49 54.48
-Appendix F Convergence Analysis
+[…]
+Table 14: Performance on low heterogeneous (Bloom-560M). Each dataset is split across two clients. See Table 5 for dataset mapping.
+Appendix G Convergence Analysis
 Since our mixer depends on each sample input, existing convergence analysis frameworks cannot be directly applied.
@@
 Nevertheless, our initial analysis provides valuable insights into the convergence behavior and offers a theoretical foundation that complements our empirical findings.
-F.1 Notation and Setup
+Notation and Setup
 Let the trainable parameters for client $k$ at round $t$ be denoted by
@@
 We denote the loss function for client $k$ as $L_{k}(Z_{k}|\mathbf{A}_{k}^{R},\mathbf{B}_{k}^{R})$, which measures the performance on client $k$’s local dataset given its trainable parameters and the fixed RoW LoRA components.
-F.2 Assumptions
-Assumption F.1 (Smoothness).
+Assumptions
+Assumption 1 (Smoothness).
 For each client $k$, the loss function $L_{k}$ is $\mu$-smooth with respect to $Z_{k}$, i.e., for any $Z_{k}^{1},Z_{k}^{2}$:
 $\|\nabla L_{k}(Z_{k}^{1})-\nabla L_{k}(Z_{k}^{2})\|\leq\mu\|Z_{k}^{1}-Z_{k}^{2}\|$ (7)
-Assumption F.2 (Strong Convexity).
+Assumption 2 (Strong Convexity).
 For each client $k$, the loss function $L_{k}$ is $\lambda$-strongly convex with respect to $Z_{k}$, i.e., for any $Z_{k}^{1},Z_{k}^{2}$:
 $L_{k}(Z_{k}^{2})\geq L_{k}(Z_{k}^{1})+\langle\nabla L_{k}(Z_{k}^{1}),Z_{k}^{2}-Z_{k}^{1}\rangle+\frac{\lambda}{2}\|Z_{k}^{2}-Z_{k}^{1}\|^{2}$ (8)
-Assumption F.3 (Bounded RoW Sensitivity).
+Assumption 3 (Bounded RoW Sensitivity).
 The optimal parameters for client $k$ are sensitive to changes in the RoW parameters with a bounded Lipschitz constant.
 Specifically, for any two sets of RoW parameters $(\mathbf{A}_{k}^{R,1},\mathbf{B}_{k}^{R,1})$ and $(\mathbf{A}_{k}^{R,2},\mathbf{B}_{k}^{R,2})$:
-$\|Z_{k}^{*}(\mathbf{A}_{k}^{R,1},\mathbf{B}_{k}^{R,1})-Z_{k}^{*}(\mathbf{A}_{k}^{R,2},\mathbf{B}_{k}^{R,2})\|\leq\beta\|(\mathbf{A}_{k}^{R,1},\mathbf{B}_{k}^{R,1})-(\mathbf{A}_{k}^{R,2},\mathbf{B}_{k}^{R,2})\|$ (9)
+$\|Z_{k}^{*}(\mathbf{A}_{k}^{R,1},\mathbf{B}_{k}^{R,1})-Z_{k}^{*}(\mathbf{A}_{k}^{R,2},\mathbf{B}_{k}^{R,2})\|$
+$\leq\beta\|(\mathbf{A}_{k}^{R,1},\mathbf{B}_{k}^{R,1})-(\mathbf{A}_{k}^{R,2},\mathbf{B}_{k}^{R,2})\|$ (9)
 where $Z_{k}^{*}(\mathbf{A}_{k}^{R},\mathbf{B}_{k}^{R})=\arg\min_{Z_{k}}L_{k}(Z_{k}|\mathbf{A}_{k}^{R},\mathbf{B}_{k}^{R})$ represents the optimal parameters for client $k$ given fixed RoW parameters.
@@
 The third assumption holds because the RoW LoRA component remains frozen during local training, providing a stable reference point that limits how much client parameters can drift between rounds based on changes in other clients’ parameters.
-F.3 Convergence Theorem
-Theorem F.4 (Convergence of FedALT).
+Convergence Theorem
+Theorem 1 (Convergence of FedALT).
 With the assumptions, we can derive:
@@
 From Assumption 1 (smoothness), we can have:
-$L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ $\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})+\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{t,s+1}-Z_{k}^{t,s}\rangle$
+$L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
+$\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
+$\quad+\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{t,s+1}-Z_{k}^{t,s}\rangle$
 $\quad+\frac{\mu}{2}\|Z_{k}^{t,s+1}-Z_{k}^{t,s}\|^{2}$ (11)
 Substituting the update rule:
-$L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ $\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
+$L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
+$\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
 $\quad+\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),-\eta\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\rangle$
-$\quad+\frac{\mu}{2}\|-\eta\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (12)
+$\quad+\frac{\mu}{2}\|-\eta\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$
 $=L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-\eta\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$
-$\quad+\frac{\mu\eta^{2}}{2}\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (13)
-$=L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-\eta(1-\frac{\mu\eta}{2})\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (14)
+$\quad+\frac{\mu\eta^{2}}{2}\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$
+$=L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
+$\quad-\eta(1-\frac{\mu\eta}{2})\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (12)
 Let $Z_{k}^{*}=Z_{k}^{*}(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ denote the optimal parameters that minimize the loss function for client $k$ given the fixed RoW parameters.
@@
 $L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
-$\geq$ $L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})+\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{*}-Z_{k}^{t,s}\rangle+\frac{\lambda}{2}\|Z_{k}^{*}-Z_{k}^{t,s}\|^{2}$ (15)
+$\geq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
+$\quad+\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{*}-Z_{k}^{t,s}\rangle$
+$\quad+\frac{\lambda}{2}\|Z_{k}^{*}-Z_{k}^{t,s}\|^{2}$ (13)
 Rearranging:
 $\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{t,s}-Z_{k}^{*}\rangle$
-$\geq$ $L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})+\frac{\lambda}{2}\|Z_{k}^{*}-Z_{k}^{t,s}\|^{2}$ (16)
+$\geq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
+$\quad+\frac{\lambda}{2}\|Z_{k}^{*}-Z_{k}^{t,s}\|^{2}$ (14)
 By Cauchy-Schwarz inequality:
-$\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|\cdot\|Z_{k}^{t,s}-Z_{k}^{*}\|\geq\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{t,s}-Z_{k}^{*}\rangle$ (17)
+$\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|\cdot\|Z_{k}^{t,s}-Z_{k}^{*}\|$
+$\geq\langle\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}),Z_{k}^{t,s}-Z_{k}^{*}\rangle$ (15)
 Combining with the previous inequality:
 $\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|\cdot\|Z_{k}^{t,s}-Z_{k}^{*}\|$
-$\geq$ $L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})+\frac{\lambda}{2}\|Z_{k}^{*}-Z_{k}^{t,s}\|^{2}$ (18)
+$\geq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
+$\quad+\frac{\lambda}{2}\|Z_{k}^{*}-Z_{k}^{t,s}\|^{2}$ (16)
 From the properties of strongly convex functions, we can derive:
-$\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}\geq 2\lambda(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}))$ (19)
+$\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$
+$\geq 2\lambda(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}))$ (17)
 Substituting this into our earlier expression for the loss after one update:
 $L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
-$\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-\eta(1-\frac{\mu\eta}{2})\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (20)
 $\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
-$-2\eta(1-\frac{\mu\eta}{2})\lambda(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}))$ (21)
-$=(1-2\eta\lambda(1-\frac{\mu\eta}{2}))(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})),$ (22)
+$\quad-\eta(1-\frac{\mu\eta}{2})\|\nabla L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|^{2}$ (18)
+$\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
+$\quad-2\eta(1-\frac{\mu\eta}{2})\lambda(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
+$\quad-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}))$ (19)
+$=(1-2\eta\lambda(1-\frac{\mu\eta}{2}))(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
+$\quad-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})),$ (20)
 where the learning rate $\eta<\frac{2}{\mu}$.
 From Assumption 2 (strong convexity), we have:
-$\frac{\lambda}{2}\|Z_{k}^{t,s}-Z_{k}^{*}\|^{2}\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ (23)
+$\frac{\lambda}{2}\|Z_{k}^{t,s}-Z_{k}^{*}\|^{2}$
+$\leq L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ (21)
 And from Assumption 1 (smoothness):
-$L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\leq\frac{\mu}{2}\|Z_{k}^{t,s}-Z_{k}^{*}\|^{2}$ (24)
+$L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
+$\leq\frac{\mu}{2}\|Z_{k}^{t,s}-Z_{k}^{*}\|^{2}$ (22)
 Combining these inequalities with our convergence rate:
-$\frac{\lambda}{2}\|Z_{k}^{t,s+1}-Z_{k}^{*}\|^{2}$ $\leq L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ (25)
-$\leq(1-2\eta\lambda(1-\frac{\mu\eta}{2}))(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}))$ (26)
+$\frac{\lambda}{2}\|Z_{k}^{t,s+1}-Z_{k}^{*}\|^{2}$
+$\leq L_{k}(Z_{k}^{t,s+1}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$ (23)
+$\leq(1-2\eta\lambda(1-\frac{\mu\eta}{2}))(L_{k}(Z_{k}^{t,s}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})$
+$\quad-L_{k}(Z_{k}^{*}|\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t}))$ (24)
 $\leq(1-2\eta\lambda(1-\frac{\mu\eta}{2}))\frac{\mu}{2}\|Z_{k}^{t,s}-Z_{k}^{*}\|^{2}$ (25)
@@
 Consider the difference between parameters at consecutive rounds:
-$\|Z_{k}^{t+1}-Z_{k}^{t}\|$ $\leq\|Z_{k}^{t+1}-Z_{k}^{*}(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|$ (32)
-$+\|Z_{k}^{*}(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-Z_{k}^{*}(\mathbf{A}_{k}^{R,t-1},\mathbf{B}_{k}^{R,t-1})\|$ (33)
-$+\|Z_{k}^{*}(\mathbf{A}_{k}^{R,t-1},\mathbf{B}_{k}^{R,t-1})-Z_{k}^{t}\|$ (34)
+$\|Z_{k}^{t+1}-Z_{k}^{t}\|$
+$\leq\|Z_{k}^{t+1}-Z_{k}^{*}(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})\|$
+$\quad+\|Z_{k}^{*}(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-Z_{k}^{*}(\mathbf{A}_{k}^{R,t-1},\mathbf{B}_{k}^{R,t-1})\|$
+$\quad+\|Z_{k}^{*}(\mathbf{A}_{k}^{R,t-1},\mathbf{B}_{k}^{R,t-1})-Z_{k}^{t}\|$ (30)
 By Step 1 and Assumption 3:
@@
 From the definition of RoW LoRA:
-$\|(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-(\mathbf{A}_{k}^{R,t-1},\mathbf{B}_{k}^{R,t-1})\|$ $=\frac{1}{K-1}\|\sum_{m\neq k}(Z_{m}^{t}-Z_{m}^{t-1})\|$ (36)
+$\|(\mathbf{A}_{k}^{R,t},\mathbf{B}_{k}^{R,t})-(\mathbf{A}_{k}^{R,t-1},\mathbf{B}_{k}^{R,t-1})\|$
+$=\frac{1}{K-1}\|\sum_{m\neq k}(Z_{m}^{t}-Z_{m}^{t-1})\|$ (32)
 $\leq\frac{1}{K-1}\sum_{m\neq k}\|Z_{m}^{t}-Z_{m}^{t-1}\|$ (33)
@@
 When this condition is satisfied, FedALT converges to a stable solution within a neighborhood determined by the local training precision.
-Appendix G Limitations of FedALT in Traditional FL
+Appendix H Limitations of FedALT in Traditional FL
 In FedALT, a separate RoW LoRA is maintained alongside each client’s Individual LoRA component.